### v2

# Double Higgs Production via Vector Boson Fusion in SMEFT

Abstract
While gluon fusion dominates Higgs pair production at the LHC, vector boson fusion (VBF) offers a unique window into Beyond the Standard Model (BSM) physics through its distinctive kinematic features and direct sensitivity to Higgs-vector boson interactions. We perform a comprehensive analysis of double Higgs production via VBF in the Standard Model Effective Field Theory (SMEFT), systematically investigating how dimension-6 and dimension-8 bosonic operators – particularly those involving field derivatives – can enhance the production rate. We identify the most relevant Wilson coefficients (WCs) affecting the trilinear Higgs coupling (hhhhhh) and Higgs-vector boson interactions ($hVV$, hhVVhhVV). Using constraints from global fits and interpolating fit results for unconstrained WCs with Naive Dimensional Analysis, we assess their effects on the $VV\to hh$ ($V=W,Z$) scattering amplitudes and cross-sections. Our analysis includes a study of EFT convergence and validity in models with scalar extensions of the SM. Numerical simulations for the planned High-Luminosity LHC experiment (HL-LHC) in general reveal only a modest and challenging to detect enhancement of the VBF di-Higgs production rate over the SM prediction. However, we show that in optimistic scenarios, such a process could be observed at the HL-LHC. In certain cases, when the enhancement is dominated by the dimension-6 or dimension-8 operators containing field derivatives (and thus leading to stronger energy-dependent effects), this channel becomes competitive with di-Higgs production via gluon fusion. This work highlights the role of VBF di-Higgs production as a complementary channel for probing anomalous Higgs couplings and their impact on BSM physics.
1 Introduction
The discovery of the Higgs boson at the Large Hadron Collider (LHC) in 2012 [1, 2] marked a pivotal moment in particle physics, confirming the mechanism of electroweak symmetry breaking (EWSB) as described by the Standard Model (SM) [3, 4, 5]. However, there are couplings present in the SM have not yet been measured experimentally. Among them, the trilinear Higgs self-coupling is perhaps one of the most important, as it is vital for determining the exact shape of the Higgs potential and for assessing the stability of the Universe. This coupling can be probed at the LHC through non-resonant Higgs boson pair production (commonly termed as di-Higgs production), which is one of the most promising directions in the search for the physics beyond the Standard Model (BSM). Existing observed and expected $95\%$ CL experimental upper limits for inclusive gluon fusion (ggF) and vector boson fusion (VBF) double Higgs production modes read (with significant progress expected with the High-Luminosity LHC upgrade):
[…]
These numbers indicate significant potential for new physics to affect this process by enhancing its cross-section relative to the SM prediction, leading lately to growing interest in studying the process of double Higgs boson production.
The two main production channels for Higgs boson pair production in the SM are gluon fusion and vector boson fusion processes [8, 9, 10, 11]. While gluon fusion $(gg\to hh)$ [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] dominates in the SM, vector boson fusion $(qq\to qqVV\to qqhh)$ [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] constitutes a complementary channel with distinct experimental signatures, including presence of forward jets and low hadronic activity in the central region. Moreover, although VBF di-Higgs production is suppressed by an order of magnitude relative to ggF channel, it is highly sensitive to anomalous Higgs couplings - namely, the trilinear coupling hhhhhh and Higgs-gauge boson couplings $hVV$ and $hhVV$ (where “VV” stands for either a $W$-boson or a $Z$-boson). Such couplings commonly arise in various BSM theories. For reference, the leading-order tree-level diagrams contributing to di-Higgs VBF production at LHC are shown in Figure 1, with anomalous couplings marked by red blobs.
Figure 1: Diagrams contributing to the double Higgs production via vector boson fusion at tree-level and in unitary gauge (the $u$-channel is not shown).
The Standard Model Effective Field Theory (SMEFT) [59, 60, 61, 62, 63] provides a model-independent framework to parameterise BSM effects through higher-dimensional operators, suppressed by a cutoff scale $\Lambda$. These operators modify Higgs and gauge boson interactions, leading to deviations in cross-sections and kinematic distributions. In the context of double Higgs VBF production, dimension-6 and dimension-8 operators can, in principle, significantly alter the scattering amplitudes, offering a window into new physics entering the trilinear Higgs coupling that may be inaccessible in single-Higgs processes.
Over the years, numerous studies have investigated double Higgs production in the context of various BSM scenarios [64, 65, 66, 67, 68, 69, 70, 71]. A number of recent analyses have employed EFT approach, including both SMEFT and Higge Effective Field Theory (HEFT) [72, 73, 74, 75, 76], to study single, double, and multi Higgs production. Some focused on the ggF channel [77, 78, 79, 80, 81, 82, 83, 84, 85], while others examined modifications to the trilinear Higgs coupling [86, 87, 88, 89, 90, 91, 92], or concentrated on the VBF channel [93, 94, 95, 96, 97, 98, 99, 100].
In this work, we investigate double Higgs production via VBF in the SMEFT framework, examining the impact of higher-dimensional operators on the production rate and differential distributions, including terms $\propto\frac{1}{\Lambda^{2}}$ (LO in the EFT expansion) and $\propto\frac{1}{\Lambda^{4}}$ (NLO in the EFT expansion). We identify the most sensitive operators and assess the prospects for detecting BSM effects at current (LHC) and future (HL-LHC) collider experiments. Our analysis highlights the importance of VBF di-Higgs production as a unique probe of Higgs interactions and the Higgs sector, and underlines the potential of SMEFT as a tool for uncovering indirect signs of new physics. Finally, utilising SMEFT as a tool should be treated with care, as we also illustrate in this study.
The paper is organised as follows. Section 2 identifies the dimension-6 and dimension-8 SMEFT operators relevant for this analysis, and outlines our assumptions leading to bounds on maximal values of their Wilson coefficients used further in numerical analysis. In Section 3 we discuss the EFT convergence and accuracy, illustrating it with examples of the three BSM models involving heavy scalar singlets and triplets, by comparing the amplitudes calculated in renormalisable UV-theories with the corresponding ones obtained in the decoupling limit. Analytical formulae for $VV\to hh$ helicity amplitudes and cross-sections are presented in Section 4 and Appendices A and B. Finally, Section 5 contains the results of our numerical simulations, directly addressing the question of maximal potential enhancement of VBF di-Higgs production in SMEFT for LHC and HL-LHC experiments. Section 6 concludes the paper.
2 Contributing operators and bounds on Wilson coefficients
2.1 Relevant operators
To affect the process of double Higgs production via VBF, the SMEFT operators can be inserted into three types of vertices: hhhhhh, hVVhVV, and $hhVV$ (red dots on diagrams in Figure 1, $h$ indicates the Higgs and $V$ indicates the gauge boson). The relevant classes of the SMEFT operators that can impact this process are all summarised in Table 1.
Operator class
Vertex Dimension-6 Dimension-8
hhhhhh $\varphi^{6}$, $\varphi^{4}D^{2}$ $\varphi^{8}$, $\varphi^{6}D^{2}$
hVVhVV $\varphi^{4}D^{2}$, $X^{2}\varphi^{2}$ $\varphi^{6}D^{2}$, $X^{2}\varphi^{4}$, $X\varphi^{4}D^{2}$
hhVVhhVV $\varphi^{4}D^{2}$, $X^{2}\varphi^{2}$ $\varphi^{4}D^{4}$, $\varphi^{6}D^{2}$, $X^{2}\varphi^{4}$, $X^{2}\varphi^{2}D^{2}$
Table 1: Classes of SMEFT dimension-6 and dimension-8 operators affecting vertices relevant for double Higgs boson production via VBF.
Table 2 presents the specific operators that may affect one or both of the two VBF processes of interest: $WW\rightarrow hh$ and $ZZ\rightarrow hh$. All these operators have been included in the corresponding matrix elements presented in Section 4 and analytical results are given in Appendices A and B.
Dimension-6 Dimension-8
$\varphi^{6}$ $\varphi^{8}$
$Q_{\varphi}$ $(\varphi^{\dagger}\varphi)^{3}$ $Q_{\varphi^{8}}$ $(\varphi^{\dagger}\varphi)^{4}$
$\varphi^{4}D^{2}$ $\varphi^{4}D^{4}$
$Q_{\varphi\Box}$ $(\varphi^{\dagger}\varphi)\raisebox{-1.42262pt}{$\Box$}(\varphi^{\dagger}\varphi)$ $Q_{\varphi^{4}D^{4}}^{(1)}$ $(D_{\mu}\varphi^{\dagger}D_{\nu}\varphi)(D^{\nu}\varphi^{\dagger}D^{\mu}\varphi)$
$Q_{\varphi D}$ $\mathopen{}\mathclose{{\left(\varphi^{\dagger}D^{\mu}\varphi}}\right)^{\ast}\mathopen{}\mathclose{{\left(\varphi^{\dagger}D_{\mu}\varphi}}\right)$ $Q_{\varphi^{4}D^{4}}^{(2)}$ $(D_{\mu}\varphi^{\dagger}D_{\nu}\varphi)(D^{\nu}\varphi^{\dagger}D^{\mu}\varphi)$
$Q_{\varphi^{4}D^{4}}^{(3)}$ $(D_{\mu}\varphi^{\dagger}D^{\mu}\varphi)(D_{\nu}\varphi^{\dagger}D^{\nu}\varphi)$
$\varphi^{6}D^{2}$
$Q_{\varphi^{6}D^{2}}$ $(\varphi^{\dagger}\varphi)(\varphi^{\dagger}D_{\mu}\varphi)^{\ast}(\varphi^{\dagger}D^{\mu}\varphi)$
$Q_{\varphi^{6}\square}$ $(\varphi^{\dagger}\varphi)^{2}\square(\varphi^{\dagger}\varphi)$
$X^{2}\varphi^{2}$ $X^{2}\varphi^{4}$
$Q_{\varphi W}$ $\varphi^{\dagger}\varphi\,W^{I}_{\mu\nu}W^{I\mu\nu}$ $Q_{W^{2}\varphi^{4}}^{(1)}$ $(\varphi^{\dagger}\varphi)^{2}W^{I}_{\mu\nu}W^{I\mu\nu}$
$Q_{\varphi B}$ $\varphi^{\dagger}\varphi\,B_{\mu\nu}B^{\mu\nu}$ $Q_{B^{2}\varphi^{4}}^{(1)}$ $(\varphi^{\dagger}\varphi)^{2}B_{\mu\nu}B^{\mu\nu}$
$Q_{\varphi WB}$ $\varphi^{\dagger}\tau^{I}\varphi\,W^{I}_{\mu\nu}B^{\mu\nu}$ $X^{2}\varphi^{2}D^{2}$
$Q_{W^{2}\varphi^{2}D^{2}}^{(1)}$ $(D^{\mu}\varphi^{{\dagger}}D^{\nu}\varphi)W_{\mu\rho}^{I}W_{\nu}^{I\rho}$
$Q_{W^{2}\varphi^{2}D^{2}}^{(2)}$ $(D^{\mu}\varphi^{\dagger}\tau^{I}D^{\nu}\varphi)W_{\mu\rho}^{I}W_{\nu}^{\rho}$
$Q_{B^{2}\varphi^{2}D^{2}}^{(1)}$ $(D^{\mu}\varphi^{{\dagger}}D^{\nu}\varphi)B_{\mu\rho}B_{\nu}^{\rho}$
$Q_{B^{2}\varphi^{2}D^{2}}^{(2)}$ $(D^{\mu}\varphi^{\dagger}D_{\mu}\varphi)B_{\nu\rho}B^{\nu\rho}$
$X\varphi^{4}D^{2}$
$Q_{W\varphi^{4}D^{2}}^{(1)}$ $(\varphi^{{\dagger}}\varphi)(D^{\mu}\varphi^{{\dagger}}\tau^{I}D^{\nu}\varphi)W_{\mu\nu}^{I}$
$Q_{B\varphi^{4}D^{2}}^{(1)}$ $(\varphi^{{\dagger}}\varphi)(D^{\mu}\varphi^{{\dagger}}D^{\nu}\varphi)B_{\mu\nu}$
Table 2: Chosen bosonic SMEFT dimension-6 and dimension-8 operators affecting the $VV\rightarrow hh$ process, and consequently, double Higgs boson production via VBF. Operators have been categorised based on their field structure: $X$ denotes the field strength tensor, $D$ the covariant derivative and $\varphi$ the Higgs doublet.
Some important comments are in place here. As we are interested in the most general features of the VBF part of the process, we neglect the impact of the SMEFT operators on the external quark lines and on further decays of the Higgs boson pair into the final states. At the dimension-6 level, the primary potential contributions may originate from the dipole operators $Q_{uW}^{11},Q_{dW}^{11}$ that are chirality-flipping, as well as the chiral operators $Q_{\varphi ud}^{11}$ and $Q_{\varphi q}^{(3)}]^{11}$. Especially the latter could amplify the longitudinal $W$-boson scattering, and as it was shown in ref. [101] for $WW\to WW$ scattering, these operators result in distinct kinematic distributions, possibly leading to the same effects in $WW\to hh$. However, they are strongly constrained by the $Z$-pole precision observables, e.g. for $Q_{\varphi q}^{(3)}]^{11}$, a recent electroweak analysis [102] shows a tight bound on its corresponding WC, of less than $0.02\,\frac{1}{\mathrm{TeV}^{2}}$. Despite that, the concern about such fermionic operators affecting studied process remains, especially at the dimension-8 level. Effects of a subset of such operators have been recently studied in Ref. [85] in the case of single Higgs production. Additionally, we neglect all CP-violating operators. We also do not consider the effects of the Renormalisation Group Equation (RGE) running of the WCs. Although it has been shown in the literature (see, e.g., [80, 103, 104, 105, 106]) that RGE effects are often non-negligible, we have verified that, in our case, they can change the values of the dimension-6 WCs by about $\mathcal{O}(20\%)$. This may slightly modify our results, but with no significant impact on the final conclusions.
2.2 Bounds on Wilson coefficients
To estimate the possible impact of SMEFT operators on double Higgs production via VBF, one must assign numerical values to the Wilson coefficients. For dimension-6 WCs, one can assume the maximum values that current fits to experimental data allow, as detailed in Refs. [107, 108, 109]. These values are displayed in Table 3.
WC 95% CL bounds, $\mathcal{O}(\Lambda^{-2})$ 95% CI bounds, $\mathcal{O}(\Lambda^{-4})$
Individual Marginalised Individual Marginalised
$C_{\varphi B}$ [-0.005, 0.062] [-0.310, 0.573] [-0.005, 0.002] $\cup$ [0.085, 0.092] [-0.103, 0.152]
$C_{\varphi W}$ [-0.018, 0.006] [-0.273, 0.707] [-0.017, 0.006] $\cup$ [0.282, 0.305] [-0.067, 0.338]
$C_{\varphi WB}$ [0.007, 0.003] [-0.525, 0.504] [-0.007, 0.003] [-0.190, 0.263]
$C_{\varphi\Box}$ [0.416, 1.193] [-1.715, 1.879] [-0.429, 1.141] [-1.856, 1.199]
$C_{\varphi D}$ [-0.027, -0.003] [-1.063, 1.149] [-0.027, -0.003] [-0.513, 0.483]
Table 3: The 95% CL bounds on the EFT coefficients from linear, $\mathcal{O}(\Lambda^{-2})$, and quadratic, $\mathcal{O}(\Lambda^{-4})$ EFT fits taken from [109]. $\Lambda=1$ TeV is assumed. Results from individual and marginalised fits are presented separately.
The maximum allowed values for $C_{\varphi}$ (not shown in Table 3) can be obtained from the current allowed 95% confidence level intervals for the trilinear Higgs coupling modifier $\kappa_{\lambda_{3}}=\lambda_{3}/\lambda_{3}^{SM}$. Bounds on $\kappa_{\lambda_{3}}$ derived from combined single- and double Higgs production searches read (with other couplings set to their SM predictions) [^4] There is also new CMS measurement constraining $\kappa_{\lambda_{3}}$ to similar range, $\kappa_{\lambda_{3}}\in[-1.39,\,7.02]$, based on di-Higgs production only [7].:
[…]
However, within linear field redefinitions in SMEFT, the trilinear Higgs coupling hhhhhh is a momentum-dependent quantity through the WCs $C_{\varphi D}$ and $C_{\varphi\Box}$ and related WCs associated with dimension-8 operators. Under the assumption that all other WCs are negligibly small compared to $C_{\varphi}$ and $C_{\varphi^{8}}$, this limit can be translated into a bound on the following combination in the $\{G_{F},m_{W},m_{Z}\}$-input parameter scheme [112, 84],
$\kappa_{\lambda_{3}}=1-\frac{C_{\varphi}}{(G_{F}m_{h}^{2})(G_{F}\Lambda^{2})}-\frac{\sqrt{2}\,C_{\varphi^{8}}}{(G_{F}m_{h}^{2})(G_{F}\Lambda^{2})^{2}}\;,$ (2.2)
where $\lambda_{3}^{SM}=\sqrt{2}G_{F}m_{h}^{2}$. Assuming $C_{\varphi^{8}}=0$ and $\Lambda=1$ TeV, one can translate the bound on $\kappa_{\lambda_{3}}$ from Eq. (2.1) into the direct bound on $C_{\varphi}$. Using the less stringent constraints from CMS one gets,
$\frac{13.8}{\text{TeV}^{2}}\lesssim C_{\varphi}\lesssim\frac{4.7}{\text{TeV}^{2}}\;.$ (2.3)
While our study adopts conservative bounds on $C_{\varphi}$, we note that recent trilinear coupling analysis [113, 114] suggests that somewhat stronger constraints may apply.
In the case of dimension-8 Wilson coefficients, there have been some attempts to impose constraints based on available experimental data and theoretical considerations, such as unitarity bounds [71, 115]. However, these constraints are significantly weaker than the values adopted in this work following assumptions, which we list below. Consequently, we do not include them in our analysis. Instead, one can roughly estimate their maximal values using the approach of [116], based on the study of the validity of the EFT expansion, which leads to a particular power-counting of WCs, also known as Naive Dimensional Analysis (NDA),
$\frac{C_{i}}{\Lambda^{D-4}}=\frac{g_{*}^{n_{i}-2}\times c_{i}}{\Lambda^{D-4}}\;,$ (2.4)
where $D$ is the dimension of SMEFT operator, $n_{i}$ is the number of fields involved, $g_{*}\leq 4\pi$ is a generic UV coupling, and $c_{i}$ is an additional suppression factor related, e.g., to additional symmetries such as shift symmetry or custodial symmetry. As a result, we get,
$C_{\varphi^{8}}$ $=g_{*}^{6}\,c_{\varphi^{8}}\;,\quad C_{\varphi^{6}\Box}=g_{*}^{4}\,c_{\varphi^{6}\Box}\;,\quad C_{\varphi^{6}D^{2}}=g_{*}^{4}\,c_{\varphi^{6}D^{2}}\;,\quad C_{\varphi^{4}D^{4}}^{(i)}=g_{*}^{2}\,c_{\varphi^{4}D^{4}}^{(i)}\;,$ (2.5)
$C_{V^{2}\varphi^{4}}^{(i)}$ $=g_{*}^{4}\,c_{V^{2}\varphi^{4}}^{(1)}\;,\qquad C_{V^{2}\varphi^{2}D^{2}}^{(i)}=g_{*}^{2}\,c_{V^{2}\varphi^{2}D^{2}}^{(1)}\;,\qquad C_{V\varphi^{4}D^{2}}^{(i)}=g_{*}^{3}\,c_{V\varphi^{4}D^{2}}^{(1)}\;\;,$
with $V=W$ or $B$.
Further assumption of the maximal possible value of $g_{*}=4\pi$ and no additional suppression factors ($c_{i}=1$) is in principle allowed in strongly coupled UV scenarios, but may lead to problems with EFT expansion convergence. Let us consider as an example the Higgs boson mass in SMEFT. At the tree level, it receives contributions from all higher order operators of the form $Q_{\varphi^{n}}(\varphi^{\dagger}\varphi)^{n+3}$. After EWSB, assuming the NDA form of (2.4) for $C_{\varphi^{n}}$, one has [117],
$m_{h}^{2}=\frac{1}{Z_{h}^{2}}\mathopen{}\mathclose{{\left[\lambda v^{2}-g_{*}^{2}v^{2}\sum_{n=0}^{\infty}(n+2)(n+3)c_{\varphi^{n}}\,\mathopen{}\mathclose{{\left(\frac{g_{*}^{2}v^{2}}{2\Lambda^{2}}}}\right)^{n+1}}}\right]\;,$ (2.6)
where the constant $Z_{h}$ is required to ensure canonical normalisation of the Higgs field in the kinetic terms, and is dependent on WCs of the form $\varphi^{4+2n}D^{2}\;,n=0,1,2,\dots$ in Table 2. Assuming that all $c_{\varphi^{n}}$ are of order ${\cal O}(1)$ (or at least not growing faster than $n$ to some fixed power), such a power series converges if,
$\frac{g_{*}^{2}v^{2}}{2\Lambda^{2}}<1\;.$ (2.7)
Incidentally, Eq. (2.6) displays a certain fine-tuning for keeping $m_{h}^{2}\ll\Lambda^{2}$, when the upper bound of (2.7) is nearly saturated. Similar constraints on WCs containing Higgs field derivatives follow from the requirements of consistency of the SMEFT Lagrangian. E.g. $Z_{h}$ factor at dimension-6 level has the form [118],
$Z_{h}=1+\frac{1}{4}C_{\varphi D}\,\frac{v^{2}}{M^{2}}-C_{\varphi\Box}\,\frac{v^{2}}{M^{2}}\,.$ (2.8)
This has to be greater than zero (otherwise a Higgs tachyon in SMEFT appears), leading to the condition,
$\frac{|4\,C_{\varphi\Box}-C_{\varphi D}|v^{2}}{\Lambda^{2}}<\sqrt{2}\;,$ (2.9)
or equivalently, using NDA for WCs through Eq. (2.5), to a bound,
$\frac{g_{*}^{2}v^{2}}{\Lambda^{2}}\lesssim{\cal O}(1)\;.$ (2.10)
Consequently, in the context of NDA estimates for the upper bounds on the magnitudes of Wilson coefficient, the assumption $g_{*}\sim 4\pi$ is generally invalid when the scale $\Lambda$ is below approximately 3.1 TeV, where to ensure the convergence of the EFT series, the effective coupling $g_{*}$ should satisfy a tighter constraint $g_{*}\lesssim\frac{\Lambda}{v}$.
Additionally, assuming the NDA scaling of Eq. (2.4), the bounds on dimension-6 Wilson coefficients obtained from fits to the experimental data set a natural limit on UV coupling $g_{*}$. The most stringent limit arises from constraints on $C_{\varphi}$ in Eq. (2.3), leading to,
$g_{*}\lesssim\mathrm{min}\mathopen{}\mathclose{{\left(|C_{\varphi}^{\mathrm{m}ax}|^{1/4}\,\sqrt{\frac{\Lambda}{\text{1 TeV}}},\,4\pi}}\right)\,$, (2.11)
where, to be conservative, we use a maximal allowed value of $C_{\varphi}^{\mathrm{m}ax}=13.8$ derived from Eq. (2.3). Likewise, applying the NDA estimation to the Wilson coefficient $C_{\varphi\Box}$ yields the following bound,
$g_{*}\lesssim\mathrm{min}\mathopen{}\mathclose{{\left(\sqrt{|C_{\varphi\Box}^{\mathrm{m}ax}|}\,\frac{\Lambda}{\text{1 TeV}},\,4\pi}}\right)\,$, (2.12)
with $|C_{\varphi\Box}^{\mathrm{m}ax}|\approx 2$ following from Table 3. For our estimates of possible enhancements to double Higgs production via the VBF channel in the SMEFT, we use the $g_{*}$ limits displayed above as the maximal NDA value for WCs that currently have no experimental constraints. As operators with and without derivatives may arise from distinct underlying physics processes, we have chosen to employ different bounding schemes for each category, using the constraint (2.11) for $C_{\varphi^{8}}$ and (2.12) for the remaining WCs not yet included in fits to experimental data.
Indeed, the above-mentioned “educated-NDA” procedure works quite well for keeping the EFT convergence under control. To illustrate this point, if we take the scaling of $C_{\varphi^{8}}$ from Eq. (2.5) and impose the bound from Eq. (2.11), we arrive at ($\Lambda$ expressed in TeV):
$\frac{C_{\varphi^{8}}}{\Lambda^{4}}\sim\frac{g_{*}^{6}}{\Lambda^{4}}\sim\frac{10^{3/2}}{\Lambda}\;.$ (2.13)
This demonstrates that $\frac{C_{\varphi^{8}}}{\Lambda^{4}}$ decreases as $\Lambda$ grows and remains bounded. Critically, assuming a large coupling $g_{*}=4\pi$ even at a relatively high scale like $\Lambda=10\,\text{TeV}$ would lead to a drastically different result, $\frac{C_{\varphi^{8}}}{\Lambda^{4}}=125\pi$, which is orders of magnitude larger and likely inconsistent with Eq. (2.2) and the convergence of EFT expansion.
In summary, to check the role of VBF processes in di-Higgs production, we employ the following bounds. For WCs corresponding to a non-derivative operator $Q_{\varphi^{8}}$, we use the experimental CMS bound [111] on $\kappa_{\lambda_{3}}$ and the resulting limits on $C_{\varphi}$ in Eq. (2.3) together with the NDA scaling of (2.4) for $c_{\varphi^{8}}=1$. We then obtain,
$-\mathrm{min}\mathopen{}\mathclose{{\left(1.8\,\sqrt{\frac{\Lambda}{\text{1 TeV}}},\,4\pi}}\right)\,\lesssim$ $\,g_{*}^{\prime}\lesssim\mathrm{min}\mathopen{}\mathclose{{\left(1.4\,\sqrt{\frac{\Lambda}{\text{1 TeV}}},\,4\pi}}\right)\,$, (2.14)
$-\frac{31.6}{\text{TeV}^{4}}$ $\lesssim\frac{C_{\varphi^{8}}}{\Lambda^{4}}\lesssim\frac{7.4}{\text{TeV}^{4}}\;.$
However, as can be seen from the analytical expressions of the helicity amplitudes $VV\to hh$ given in Appendices A and B, the dominant contribution to di-Higgs production both in $WW$ or $ZZ$ scattering is proportional to the same combination of $C_{\varphi}$ and $C_{\varphi^{8}}$ as the one entering $\kappa_{\lambda_{3}}$ in Eq. (2.2). For this reason, individual estimates of both WCs do not significantly affect numerical results and remain bounded primarily by already known experimental limits on $\kappa_{\lambda_{3}}$, as we discuss further in Section 5.
For derivative operators, again being conservative, we assume maximal values slightly larger than those derived from global fits for marginalised and linear case presented in Table 3,
$\mathopen{}\mathclose{{\left|C_{\varphi\Box}}}\right|\lesssim\frac{2}{\text{TeV}^{2}},\quad\mathopen{}\mathclose{{\left|C_{\varphi D}}}\right|\lesssim\frac{1.5}{\text{TeV}^{2}},\quad\mathopen{}\mathclose{{\left|C_{\varphi\mathopen{}\mathclose{{\left(W,B,WB}}\right)}}}\right|\lesssim\frac{1}{\text{TeV}^{2}}\;,$ (2.15)
which, after making use of (2.12) result in bounds,
$g_{*}$ $\lesssim\text{max}\mathopen{}\mathclose{{\left(\sqrt{2}\frac{\Lambda}{1\text{ TeV}},\,4\pi}}\right)\;,$ (2.16)
$\frac{|C_{\varphi^{6}D^{2}}|}{\Lambda^{4}}$ $=\frac{|C_{V^{2}\varphi^{4}}|}{\Lambda^{4}}=\frac{g_{*}^{4}}{\Lambda^{4}}\lesssim\frac{4}{(1\text{ TeV})^{4}}\;,$
$\frac{|C_{\varphi^{4}D^{4}}|}{\Lambda^{4}}$ $=\frac{|C_{V^{2}\varphi^{2}D^{2}}|}{\Lambda^{4}}=\frac{g_{*}^{2}}{\Lambda^{4}}\lesssim\frac{2}{(1\text{ TeV})^{2}}\>\frac{1}{\Lambda^{2}}\;,$
$\frac{|C_{W\varphi^{4}D^{2}}|}{\Lambda^{4}}$ $=\frac{|C_{V\varphi^{4}D^{2}}|}{\Lambda^{4}}=\frac{g_{*}^{3}}{\Lambda^{4}}\lesssim\frac{2\sqrt{2}}{(1\text{ TeV})^{3}}\>\frac{1}{\Lambda}\;.$
The bounds from Eqs. (2.14) and (2.16) will serve as the primary input values for the subsequent analysis.
3 Wilson coefficients and EFT convergence in scalar extensions of the SM
Several questions arise when discussing the maximal enhancement of the double Higgs boson production within SMEFT. First, are the NDA estimates realistic, i.e. can one explicitly construct UV models which after decoupling of heavy particles lead to large values of WCs? Second, if WCs are large, can they affect the convergence of the EFT expansion in inverse powers of heavy mass scale? Third, how reliable is the SMEFT approximation for the growing energy of the process?
We address these questions by considering examples of a few simple extensions of the SM. These UV-completions are models with a heavy real scalar singlet field $S$ or two types of heavy scalar triplets $T_{1}$, $T_{2}$. They provide us with a useful theoretical laboratory for studying and testing the VBF di-Higgs production in the SMEFT.
3.1 Real scalar model
The matching onto the SMEFT of a heavy scalar singlet model has been worked out in several studies, with dimension-6 operators at tree-level [119, 120] and one-loop [121, 122] as well as dimension-8 operators [123]. By incorporating all renormalisable interactions between the heavy singlet $S$ and the SM fields, we have
$\mathcal{L}_{S}=\mathcal{L}_{SM}+\frac{1}{2}(\partial_{\mu}S)^{2}-\frac{1}{2}M^{2}S^{2}-\alpha\,M\,S\varphi^{\dagger}\varphi-\frac{1}{2}\kappa\,|\varphi|^{2}S^{2}-\frac{1}{3!}\beta\,M\,S^{3}-\frac{1}{4!}\lambda_{S}\,S^{4}\;,$ (3.1)
where $\mathcal{L}_{SM}$ is the SM-Lagrangian in the notation of Refs. [63, 118] and $\varphi$ is the SM Higgs doublet. Besides the heavy singlet mass $M$, all other parameters are dimensionless. By promoting the parameters of the theory into auxiliary fields, a background $Z_{2}$-symmetry under which $\varphi\to\varphi$ and $S\to-S$, assigns charges (1)(-1) to $\alpha$ and $\beta$ and $(+1)$ to all other couplings. The Wilson coefficients in front of the $Z_{2}$-invariant EFT expansions in terms of $v^{2}/M^{2}$ (or $s/M^{2}$ when calculating momentum-dependent terms in scattering amplitudes) should also have an $Z_{2}$-invariant polynomial form,
$C_{i}=f(\alpha^{n_{1}}\beta^{n_{2}},...;,\kappa,\lambda_{S})\;,\quad n_{1}+n_{2}=\mathrm{even}\;,$ (3.2)
Alternatively to the standard matching procedure, one can try to identify SMEFT WCs by comparing the corresponding terms of the amplitudes in both UV and effective theory in the energy region $v^{2}\ll s\ll M^{2}$. Relevant to the di-Higgs production processes, we can use longitudinally polarised vector bosons $W_{L}W_{L}\to hh$ and $Z_{L}Z_{L}\to hh$ scattering, although the method works as well for other choices (e.g. $W_{L}W_{L}\to W_{L}W_{L}$ and $W_{L}W_{L}\to Z_{L}Z_{L}$). Let us compare the leading (not suppressed by additional $v^{2}/M^{2}$ powers) contributions to coefficients of $s$ powers in the amplitude. In the high $s\gg v^{2}$ regime, the Goldstone Boson Equivalence Theorem [124, 125, 126] simplifies the calculation since at high energy in SMEFT we only have to calculate $G^{+}G^{-}\to hh$ and $G^{0}G^{0}\to hh$ and match to the same amplitudes in full theory at $s\ll M^{2}$. The tree level longitudinal amplitudes in the UV model for both processes are given by the same expression, including contributions from the SM contact 4-scalar interaction and the single $s$-channel exchange of the heavy scalar (see Fig 2),
$\mathcal{M}_{00}^{(G^{+}G^{-}\to hh)}(s)=\mathcal{M}_{00}^{(G^{0}G^{0}\to hh)}(s)=-\lambda+\frac{(-i\alpha M)^{2}}{s-M^{2}}\overset{s\ll M^{2}}{=}-\lambda+\alpha^{2}\,\mathopen{}\mathclose{{\left[1+\frac{s}{M^{2}}+\frac{s^{2}}{M^{4}}\dots+}}\right]\;.$ (3.3)
where $\lambda$ is the quartic Higgs coupling in $\mathcal{L}_{SM}$: $\frac{\lambda}{2}(\varphi^{\dagger}\varphi)^{2}$.
Figure 2: Dominant diagrams contributing to the di-Higgs production in the real scalar SM extension. The black dot indicates the quartic Higgs coupling $\lambda$. The fields $G^{0},G^{\pm}$ are the Goldstone bosons in the $R_{\xi}$-gauge we are using and $h$ is the Higgs field.
It is crucial to note the same prefactor $\alpha^{2}$ for coefficients of $(s/M^{2})^{n}\;,n=0,1,2,..$ powers. On the other hand, the SMEFT amplitudes at large-$s$, but always below the singlet mass squared, give (in dimension-6 basis of [63] and dimension-8 basis of [127]; the explicit expressions for the amplitudes are presented in Appendices A and B):
$\mathcal{M}_{00}^{(G^{+}G^{-}\to hh)}(v^{2}\ll s\ll M^{2})=-\lambda_{eff}$ $-\frac{1}{2}(4C^{\varphi\Box}-C^{\varphi D})\frac{s}{M^{2}}$
$+\frac{1}{8}\biggl[(C^{(1)}_{\varphi^{4}D^{4}}+C^{(2)}_{\varphi^{4}D^{4}})(1+\cos^{2}\theta)+4\,C^{(3)}_{\varphi^{4}D^{4}}\biggr]\frac{s^{2}}{M^{4}}\;,$ (3.4)
$\mathcal{M}_{00}^{(G^{0}G^{0}\to hh)}(v^{2}\ll s\ll M^{2})=-\lambda_{eff}$ $-(2C^{\varphi\Box}+C^{\varphi D})\frac{s}{M^{2}}$
$+\frac{1}{4}\biggl[2(C^{(1)}_{\varphi^{4}D^{4}}+C^{(3)}_{\varphi^{4}D^{4}})-C^{(2)}_{\varphi^{4}D^{4}}(1-\cos^{2}\theta)\biggr]\frac{s^{2}}{M^{4}}\;,$ (3.5)
where $\theta$ is the scattering angle. By comparing the $s^{0},s^{1},s^{2}$ pieces of Eqs. (3.3), (3.4) and (3.5) we get
$\lambda_{eff}=\lambda-\alpha^{2}\;,$ (3.6)
$C_{\varphi\Box}=-\frac{\alpha^{2}}{2}\;,\quad C_{\varphi D}=0\;,$ (3.7)
$C_{\varphi^{4}D^{4}}^{(3)}=2\alpha^{2}\;,\quad C_{\varphi^{4}D^{4}}^{(1)}=C_{\varphi^{4}D^{4}}^{(2)}=0\;,$ (3.8)
where $\lambda_{eff}$ corresponds to low-energy quartic SM Higgs coupling.
These results are in agreement with the literature [119, 121, 122, 120, 123]. Note also that the dimension-8 WCs are positive semi-definite according to the corresponding positivity bound [128]. Furthermore, $C_{\varphi}$ can be found by comparing the corresponding amplitudes for $\varphi^{6}$ processes in full and effective theory. It is [122]
$C_{\varphi}=-\frac{\alpha^{2}}{2}\biggl(\kappa-\frac{\alpha\beta}{3}\biggr)\;.$ (3.9)
The above procedure suggests the matching conditions to be taken at $v^{2}\ll s=\Lambda_{0}^{2}\ll M^{2}$, with $\Lambda_{0}$ could naturally be chosen at the scale defined by the unitarity violation of SMEFT. As usual, WCs at low scale can be then calculated using the RGEs evolution.
From the UV model’s perspective, $\alpha$ is a free parameter, eventually constrained by perturbativity of couplings. In such case, the UV loop-expansion parameter should be limited to
$\frac{\alpha^{2}}{(4\pi)^{2}}\lesssim 1\;,$ (3.10)
so that $\alpha\lesssim 4\pi$ and maximal value of $C_{\varphi\Box}$ can saturate the NDA limit. Therefore, the SM extension with a real scalar provides a simple, explicit example of a model that can really produce large values for WCs in the low-energy limit [^5] However, one should note that some WCs may be constrained by additional conditions, such as the properties of the Higgs potential. Higher order operators contribute to it, ultimately affecting its stability and positions of minima. We do not study these issues in detail here, as our goal is simply to provide an example of consistent UV models which may produce large WCs..
Actually, it is possible to get even larger WCs from perturbative UV models. If instead of one heavy real scalar we add to the SM $n$ copies of such fields with couplings $\alpha_{1},\alpha_{2},\ldots$ and masses $M_{1}\leq M_{2}\leq\ldots$ (in addition we assume for simplicity that heavy scalars interact feebly with each other and mainly couple to the SM Higgs-doublet), the formula for $C_{\varphi\Box}$ takes the form,
$C_{\varphi\Box}=-\frac{1}{2M_{1}^{2}}\sum_{i=1}^{n}\alpha_{i}^{2}M_{i}^{2}\;.$ (3.11)
Obviously by increasing the number of heavy fields with similar masses one can make $C_{\varphi\Box}$ considerable, while keeping all the UV couplings $\alpha_{i}$ being perturbative.
To address the question of EFT convergence, let us note that comparison of Eq. (3.3) with Eqs. (3.4) and (3.5) shows that WCs of all higher order operators contributing to the dominant terms in the amplitude are proportional to $\alpha^{2}$ and not to the higher powers of $\alpha$. Consequently, they do not grow uncontrollably with increasing mass dimension and do not endanger the EFT series convergence. This remark is also in agreement with NDA as on dimensional ground all such couplings should be proportional to $g_{*}^{2}$. From Eq. (3.3), and, at tree level in $\hbar$-expansion, it is clear that EFT series always converges if we are below the heavy mass threshold, $s<M^{2}$, independently of WC values. Other WCs which may contribute to sub-dominant terms in the amplitude are proportional to higher powers of UV parameters and thus can be larger, but they are always multiplied by powers of $(v/M)$ and therefore should not affect EFT convergence as long as condition (2.7) is fulfilled.
3.2 Triplet scalar models
As can be seen from Eq. (3.3), the dominant contribution to $W_{L}W_{L}\to hh$ and $Z_{L}Z_{L}\to hh$ in real scalar model has always opposite sign to SM term $\lambda_{eff}$ ($\lambda_{eff}$ should be positive for Higgs potential to be stable and to have proper minima). Thus, in such a model both processes are suppressed compared to the SM, unless the $\alpha$-parameter is huge and BSM terms dominate the SM amplitude. This may not be the case for scalar triplet extensions.
Let us consider models containing heavy scalar triplets, first neutral $T_{1}(1,3,0)$ and then hypercharged $T_{2}(1,3,1)$ (numbers is parenthesis denote as usual $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$ quantum numbers). In the first case, the interesting part of the Lagrangian is
$\mathcal{L}_{T_{1}}=\mathcal{L}_{SM}+(D_{\mu}T_{1})^{\dagger}(D^{\mu}T_{1})-\alpha_{T_{1}}\>M\,\varphi^{\dagger}\sigma^{a}T_{1}^{a}\varphi+\dots$ (3.12)
where $\sigma^{a}$ are the Pauli matrices. By exploiting the same $Z_{2}$-symmetry arguments as for the real singlet scalar, we deduce that WCs will only depend on $\alpha_{T_{1}}^{2}$.
In this case, the amplitudes $\mathcal{M}_{00}^{G^{0}G^{0}\to hh}$ and $\mathcal{M}_{00}^{G^{+}G^{-}\to hh}$ are different, at high energy the first one comes from the $s$-channel only with neutral component of the triplet mediator $T_{1}^{0}$ plus the contact SM-like graph in Figure 2, while the second one also contains contributions from the $t$ and $u$ channels with charged mediators exchange (not shown in Figure 2). The amplitudes are
$\mathcal{M}_{00}^{G^{0}G^{0}\to hh}(s)=-\lambda+\frac{(i\alpha_{T_{1}}M)^{2}}{s-M^{2}}=-\lambda+\alpha_{T_{1}}^{2}\biggl[1+\frac{s}{M^{2}}+\frac{s^{2}}{M^{4}}+\dots\biggr]\;,$ (3.13)
and
$\mathcal{M}_{00}^{G^{+}G^{-}\to hh}(s)$ $=-\lambda+\alpha_{T_{1}}^{2}\>M^{2}\,\biggl[\frac{1}{s-M^{2}}-\frac{1}{t-M^{2}}-\frac{1}{u-M^{2}}\biggr]$
$=-\lambda+\alpha_{T_{1}}^{2}-2\alpha_{T_{1}}^{2}\frac{s}{M^{2}}-\frac{\alpha_{T_{1}}^{2}}{2}\frac{s^{2}}{M^{4}}(1-\cos^{2}\theta)+\dots\;.$ (3.14)
By performing the UV and EFT matching procedure we described earlier for the amplitudes in the real scalar scenario, and using the leading terms of amplitudes given in Appendices A and B, we arrive at,
$\lambda_{eff}$ $=\lambda-\alpha_{T_{1}}^{2}\;,$ (3.15)
$C_{\varphi\Box}$ $=\frac{1}{2}\alpha_{T_{1}}^{2}\;,\quad C_{\varphi D}=-2\alpha_{T_{1}}^{2}\;,$ (3.16)
$C_{\varphi^{4}D^{4}}^{(1)}$ $=4\alpha_{T_{1}}^{2}\;,\quad C_{\varphi^{4}D^{4}}^{(2)}=0\;,\quad C_{\varphi^{4}D^{4}}^{(3)}=-2\alpha_{T_{1}}^{2}\;.$ (3.17)
These results are consistent with Ref. [123] derived with functional matching. In such a model, $C_{\varphi\Box}$ is positive but non-vanishing negative $C_{\varphi D}$ is generated, so that effectively $Z_{L}Z_{L}\to hh$ cross-section is still suppressed compared to SM but $W_{L}W_{L}\to hh$ becomes enhanced. Nevertheless, both effects cannot be large due to constraints on the $C_{\varphi D}$ which enters alone in the electroweak $\rho$-parameter as [118]
$\rho-1=\frac{1}{2}\frac{v^{2}}{\Lambda^{2}}\,C_{\varphi D}\;.$ (3.18)
The $\rho$-parameter deviation from unity is constrained by the electroweak precision observables to be at one per mile level. Hence, $|\alpha_{T_{1}}|\lesssim 0.01$ and the leading neutral scalar triplet contribution affecting di-Higgs production at high energies is strongly constrained.
The situation is different in charged heavy triplet model $T_{2}(1,3,1)$, with the Lagrangian
$\mathcal{L}_{T_{2}}=\mathcal{L}_{SM}+(D_{\mu}T_{2})^{\dagger}(D^{\mu}T_{2})-\alpha_{T_{2}}\>M\,\widetilde{\varphi}^{\dagger}\sigma^{a}T_{2}^{a}\varphi\ +\ \mathrm{h.c.}\ +\dots\;.$ (3.19)
In this case, high energy scattering amplitudes are given by
$\mathcal{M}_{00}^{G^{0}G^{0}\to hh}(s)$ $=-\lambda\ +\ 2\,\alpha_{T_{2}}^{2}\,M^{2}\,\biggl[\frac{1}{s-M^{2}}-\frac{1}{t-M^{2}}-\frac{1}{u-M^{2}}\biggr]$
$=-\lambda+2\alpha_{T_{2}}^{2}-4\,\alpha_{T_{2}}^{2}\,\frac{s}{M^{2}}-\alpha_{T_{2}}^{2}(1-\cos^{2}\theta)\frac{s^{2}}{M^{4}}+\dots\;,$ (3.20)
$\mathcal{M}_{00}^{G^{+}G^{-}\to hh}(s)$ $=-\lambda\ -\alpha_{T_{2}}^{2}\,M^{2}\,\biggl[\frac{1}{t-M^{2}}+\frac{1}{u-M^{2}}\biggr]$
$=-\lambda+2\,\alpha_{T_{2}}^{2}-\alpha_{T_{2}}^{2}\,\frac{s}{M^{2}}+\frac{\alpha_{T_{2}}^{2}}{2}\,(1+\cos^{2}\theta)\,\frac{s^{2}}{M^{4}}\ +\ \dots$ (3.21)
and matching to SMEFT as before provides
$\lambda_{eff}$ $=\lambda-2\,\alpha_{T_{2}}^{2}\;,$ (3.22)
$C_{\varphi\Box}$ $=\alpha_{T_{2}}^{2}\;,\quad C_{\varphi D}=2\alpha_{T_{2}}^{2}\;,$ (3.23)
$C_{\varphi^{4}D^{4}}^{(1)}$ $=0\;,\quad C_{\varphi^{4}D^{4}}^{(2)}=4\alpha_{T_{2}}^{2}\;,\quad C_{\varphi^{4}D^{4}}^{(3)}=0\;.$ (3.24)
Comparing to the singlet and the neutral triplet scenarios, both $C_{\varphi\Box}$ and $C_{\varphi D}$ are now positive and both $Z_{L}Z_{L}\to hh$ and $W_{L}W_{L}\to hh$ channels of di-Higgs production are enhanced in this case. In addition, adding both types of triplets (with identical masses) to the SM, one has
$\lambda_{eff}$ $=\lambda-\,\alpha_{T_{1}}^{2}-2\,\alpha_{T_{2}}^{2}\;,$
$C_{\varphi\Box}$ $=\frac{1}{2}\alpha_{T_{2}}^{2}+\alpha_{T_{2}}^{2}\;,$ (3.25)
$C_{\varphi D}$ $=-2(\alpha_{T_{1}}^{2}-\alpha_{T_{2}}^{2})\;,$
$C_{\varphi^{4}D^{4}}^{(1)}$ $=4\alpha_{T_{1}}^{2}\;,\quad C_{\varphi^{4}D^{4}}^{(2)}=4\alpha_{T_{2}}^{2}\;,\quad C_{\varphi^{4}D^{4}}^{(3)}=-2\alpha_{T_{1}}^{2}\;.$ (3.26)
and in case of approximately (or exactly, due to some underlying symmetry) equal couplings, $\alpha_{T_{1}}\approx\alpha_{T_{2}}$, it is possible to get enhancement of the double Higgs production avoiding the constraints from the $\rho$-parameter. One should note that RGEs evolution can mix various WCs after decoupling and in principle lead, at lower energies, to problems with $\rho$-parameter constraints even if they were initially fulfilled. However, we have checked that in SMEFT for considered low $\Lambda$ scales such effect is small and does not affect our conclusions.
In summary, we demonstrated with explicit examples that it is possible to find renormalisable UV-models which produce at low energies large WCs in SMEFT, that may increase or decrease the di-Higgs production rates through VBF and are consistent with EFT expansion convergence.
3.3 EFT accuracy for large $\sqrt{s}$
The issue related to the EFT convergence is how well the EFT approximation reproduces the full UV-model, depending on the value of $\sqrt{s}$. For a rough estimate, in Figure 3 we plotted the relative accuracy of EFT approximation defined as
$\Delta_{EFT}=\mathopen{}\mathclose{{\left|\frac{\mathcal{M}_{00~\mathrm{(EFT)}}^{(G^{+}G^{-}\to hh)}}{\mathcal{M}_{00~\mathrm{(UV)}}^{(G^{+}G^{-}\to hh)}}}}\right|^{2}-1\;,$ (3.27)
with $\mathcal{M}_{00~\mathrm{(EFT)}}^{(G^{+}G^{-}\to hh)}$ and $\mathcal{M}_{00~\mathrm{(UV)}}^{(G^{+}G^{-}\to hh)}$ given by expressions for the real scalar or triplet models given above. Figure 3 presents results for two scenarios: the upper panels correspond to the real scalar model, and the lower panels to a Standard Model extension incorporating both neutral and charged triplets with degenerate masses and identical couplings, denoted by $\alpha_{T_{1}}=\alpha_{T_{2}}$. We also assumed that $\lambda_{eff}$ is predominantly determined by the Higgs mass value, so that $\lambda_{eff}\approx 0.26$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: The plots compare the relative accuracy of the $W_{L}W_{L}\to hh$ high energy scattering cross-section as predicted by the SMEFT and underlying UV models. The dashed lines represent $\Delta_{EFT}$ from Eq. (3.27) as a function of $\sqrt{s}/M$. The orange curve shows the EFT result with dimension-6 terms in the amplitude alone, with higher-order terms truncated in the cross-section. The blue and red curves include dimension-8 terms in the amplitude, either truncated or fully kept in its square, respectively. The dotted lines mark the $\pm 10\%$ range of relative accuracy. The solid black line shows the deviation of the cross-section ratios from unity in the scalar models relative to the SM. The upper panels present the findings for a real scalar model, and the lower panels show the findings for a model extended with both neutral and charged triplets.
Figure 3 indicates that for small couplings $\alpha(\alpha_{T_{i}})=0.2$, the EFT remains a reliable tool for all considered UV-models up to at least $\sqrt{s}/M\sim 0.7$, provided dimension-8 terms are accounted for. With large couplings, around $\mathcal{O}(1)$, the EFT approximation maintains good agreement with UV-models in scenarios with enhanced cross-sections. However, its validity is reduced to $\sqrt{s}/M\sim 0.4$ when SM-BSM interference is negative, potentially leading to a complete cancellation of the amplitude. This effect is exacerbated by the smallness of UV-amplitude in the denominator of Eq. (3.27), increasing the discrepancy. Considering the limited experimental interest in suppressed cross-sections, the EFT offers a safe approach for studying enhanced di-Higgs VBF production at the LHC, even provided that the new physics scale is not much larger than the highest investigated invariant mass of the Higgs pair. Based on the preceding analysis, we adopt the bound on the ratio of the maximal center-of-mass energy to the new physics (NP) scale to ensure EFT convergence:
$\frac{\sqrt{s_{max}}}{\Lambda}\leq x\approx 1\;.$ (3.28)
Although the value $x\approx 0.7$ is theoretically preferable for inequality (3.28), we adopt the more relaxed value $x=1$. This choice is justified by our numerical analysis in Section 5, which shows that even $x=1$ severely limits potential process enhancements. Consequently, the stricter bound $x\approx 0.7$ would yield even less favorable predictions.
It is also important to note that retaining only linear terms in dimension-6 Wilson coefficients yields considerably less accurate results. Furthermore, Figure 3 demonstrates that when the discrepancy between EFT and UV predictions becomes significant for stronger couplings. In some cases, EFT more accurately reproduces the full result if the amplitude is truncated at dimension-8 terms, even when formally higher-order terms (up to dimension-12) arising from squaring the amplitude are retained in the cross-section. More details are provided in the following subsection.
Finally, constraints on the maximum $\sqrt{s}$ for consistent EFT application can also arise from perturbative unitarity requirements (see, e.g. Refs. [129, 130, 131, 132, 101, 133, 134, 115]). However, for center-of-mass energies achievable at the LHC, unitarity violation typically does not impose stringent bounds. For instance, for WCs of operators involving Higgs doublet derivatives, it leads to a condition of the form $C\frac{s}{\Lambda^{2}}\lesssim{\cal O}(4\pi)$. This condition is generally satisfied if the relevant dimension-6 WCs remain within the bounds derived from fits to experimental data, as summarised in Eqs. (2.15) and (2.16). Similarly, our checks indicate that unitarity constraints from $2\to 3$ and $2\to 4$ processes, as discussed in Ref. [134], also do not provide meaningful constraints for $\sqrt{s}\lesssim 2$ TeV. Therefore, we do not consider such constraints in our analysis [^6] We also note that the existing literature on unitarity constraints in SMEFT generally does not consistently account for dimension-62 contributions and thus cannot be directly applied to our analysis, which includes such terms..
3.4 A note on truncation: amplitude vs cross-section
The general structure of the EFT amplitude includes terms at the leading order $\propto\frac{1}{\Lambda^{2}}$ (LO) and next-to-leading order $\propto\frac{1}{\Lambda^{4}}$ (NLO) in the EFT expansion,
$\mathcal{M}=\mathcal{M}_{\text{SM}}+\frac{1}{\Lambda^{2}}\mathcal{M}_{\text{dim6}}+\frac{1}{\Lambda^{4}}\mathcal{M}_{\text{dim6}^{2}}+\frac{1}{\Lambda^{4}}\mathcal{M}_{\text{dim8}},$ (3.29)
where $\mathcal{M}_{\text{dim6}}$ contains single insertions of dimension-6 WCs, $\mathcal{M}_{\text{dim6}^{2}}$ their double insertions, and $\mathcal{M}_{\text{dim8}}$ single insertions of dimension-8 WCs. At the level of cross-section, this leads to corrections of order as high as $\propto\frac{1}{\Lambda^{8}}$,
$\sigma\propto\mathopen{}\mathclose{{\left|\mathcal{M}}}\right|^{2}\equiv\sigma_{(\text{SM}+\text{dim6}+\text{dim6}^{2}+\text{dim8})\times(\text{SM}+\text{dim6}+\text{dim6}^{2}+\text{dim8})}\supset\frac{1}{\Lambda^{8}}\mathopen{}\mathclose{{\left|\mathcal{M}_{\text{dim6}^{2}}+\mathcal{M}_{\text{dim8}}}}\right|^{2}.$ (3.30)
In this work, we truncate only the amplitude in the EFT expansion. This means we calculate the amplitude up to a specific order but retain all higher-order terms, as those in Eq. (3.30), that arise when squaring it to obtain the cross-section. Truncation at the cross-section level would omit terms like those presented in Eq. (3.30), which are formally of dimension-12. While Figure 3 shows that amplitude truncation does not always guarantee better accuracy than cross-section truncation at a given order, a significant drawback of cross-section truncation is its potential to yield negative cross-section values. This behaviour is illustrated in Figure 4.
As a concrete example, let’s discuss the issue of truncation within the real scalar singlet model of this section. At a very high energy, the amplitude (3.3) governs the dynamics of the longitudinal vector bosons scattering $V_{L}V_{L}\to hh$ with $V=Z,W^{\pm}$,
$\mathcal{M}_{00}^{V_{L}V_{L}\to hh}(s)\ =-\lambda_{eff}\ +\ \alpha^{2}\,\biggl(\frac{s}{M^{2}}+\frac{s^{2}}{M^{4}}+\dots\biggr)\;,$ (3.31)
where $\lambda_{eff}=\lambda-\alpha^{2}$. Now we square (3.31), to find
$|\mathcal{M}_{00}^{V_{L}V_{L}\to hh}(s)|^{2}\$$=\overbrace{\underbrace{\lambda_{eff}^{2}\ -\ 2\,\lambda_{eff}\,\alpha^{2}\frac{s}{M^{2}}}_{\sigma_{1/\Lambda^{2}}}\ +\ \alpha^{4}\>\frac{s^{2}}{M^{4}}}^{|\mathcal{M}_{1/\Lambda^{2}}|^{2}}\ \ \ \overbrace{\underbrace{-2\,\lambda_{eff}\,\alpha^{2}\,\frac{s^{2}}{M^{4}}}_{\sigma_{1/\Lambda^{4}}}\ +\ 2\,\alpha^{2}\,\frac{s^{3}}{M^{6}}\ +\ \alpha^{4}\,\frac{s^{4}}{M^{8}}}^{|\mathcal{M}_{1/\Lambda^{4}}|^{2}}\;.$ (3.32)
Clearly, truncating the cross-section at order $O(\sigma_{1/\Lambda^{2}})$ or $O(\sigma_{1/\Lambda^{4}})$ can lead to negative values. In the strong coupling regime $(\alpha\gtrsim 1)$, this readily occurs for energies above $\sqrt{s}>\frac{M}{\alpha}\sqrt{\frac{\lambda_{eff}}{2}}$. Conversely, truncating the amplitude in such cases allows the non-interference term to dominate, preventing negative cross-sections. Therefore, unless the coupling $\alpha$ is weak, comparable to $\lambda_{eff}$, truncating the full amplitude is the correct approach. Issues related to truncation and EFT validity have been discussed in Refs. [83, 135, 136], where similar conclusions were reached.
Refer to caption Refer to caption
Figure 4: Comparison of truncation schemes using the ratio $R^{WW}$ between angular-integrated cross-sections in the SMEFT and the SM [c.f. Eq. (4.7)] for $C_{\varphi\Box}$ dimension-6 WC and $C_{\varphi^{6}\Box}$ dimension-8 WC (with numerical values displayed in the corresponding plots). Solid lines indicate a cross-section truncated to terms of the same maximal dimension as included in amplitude, dashed lines indicate a cross-section calculated by squaring amplitudes without any truncation.
4 Helicity amplitudes and cross-sections
Double Higgs boson production via VBF is a complicated multi-particle process in the SM, which becomes even more involved in SMEFT. For this reason, before performing the exact numerical calculations of the cross-section for the $pp\to hhjj$ process, it is instructive to study the anatomy of individual on-shell sub-amplitudes contributing to the full amplitude. More specifically, we study the on-shell scattering of opposite-sign $W^{\pm}$ bosons or a pair of neutral $Z$-bosons into a pair of Higgs bosons $h$, namely:
$V(p_{1},\lambda_{1})+V(p_{2},\lambda_{1})\rightarrow h(p_{3})+h(p_{4}),$ (4.1)
where $V=W^{\pm}$ or $Z$, $p_{i}$ are the four-momenta of the initial and final states, and $\lambda_{i}$ represent the polarizations of the initial heavy vector bosons (two transverse — $\pm$ and longitudinal — 0). The helicity amplitude for the $VV\rightarrow hh$ process can be expressed as:
$\mathcal{M}^{VV}_{\lambda_{1}\lambda_{2}}\equiv\mathcal{M}_{\lambda_{1}\lambda_{2}}^{VVHH}(\theta,s,G_{F},M_{W},M_{Z},M_{H},C_{i})\;,$ (4.2)
where $\theta\in[0,\pi]$ is the scattering angle between the incoming and outgoing particles in the Center of Mass frame.
Not all the 9 possible helicity configurations are independent. As we neglect CP-violating WCs, the following relations hold:
$\mathcal{M}^{VV}_{++}=\mathcal{M}^{VV}_{--},\quad\mathcal{M}^{VV}_{+-}=\mathcal{M}^{VV}_{-+},\quad\mathcal{M}^{VV}_{0\pm}=\mathcal{M}^{VV}_{\pm 0},$ (4.3)
leaving a total of 4 independent helicity structures. The differential cross-section for a given helicity is given by,
$\biggl(\frac{d\sigma^{VV}}{d\Omega}\biggr)_{\lambda_{1}\lambda_{2}}=\frac{1}{2!}\,\frac{1}{9}\,\frac{1}{64\pi^{2}s}\>|\mathcal{M}^{VV}_{\lambda_{1}\lambda_{2}}(\theta,s,G_{F},M_{W},M_{Z},M_{H},C_{i})|^{2}\;,$ (4.4)
and the helicity cross-section,
$\sigma(s)^{VV}_{\lambda_{1}\lambda_{2}}=\int_{0}^{2\pi}d\varphi\int_{\theta_{cut}}^{\pi-\theta_{cut}}d\theta\sin\theta\>\biggl(\frac{d\sigma^{VV}}{d\Omega}\biggr)_{\lambda_{1}\lambda_{2}}\;.$ (4.5)
By summing over all relevant helicities, the total cross-section is:
$\sigma(s)^{VV}=\sigma(s)^{VV}_{00}+2\sigma(s)^{VV}_{++}+2\sigma(s)^{VV}_{+-}+4\sigma(s)^{VV}_{0+}\;.$ (4.6)
It is important to note that, as discussed in Sec. 3, although we consistently truncate all amplitudes to dimension-8 level, we do not truncate expressions for cross-sections in the same way, keeping instead all terms arising from squaring the amplitudes. This means that we include in cross section calculations also the subset of terms up to dimension-12 (such subset is gauge-invariant [^7] We have analytically and explicitly checked (with the help of SmeftFR-code [112]) that the amplitude $VV\to hh$ is gauge invariant, i.e., gauge-fixing $\xi$-parameter independent, at tree level for dim6 and dim62 contributions alone. We expect the same holds for the dim8 amplitude terms, although we have not verified the latter cases.). For large values of WCs it can lead to numerically significant differences comparing to cross-sections truncated to dimension-8 terms only.
Explicit formulae for the leading order helicity amplitudes are provided in Appendix A for the $WW\rightarrow HH$ process and in Appendix B for the $ZZ\rightarrow HH$ process. All amplitudes and corresponding numerical results in this section were generated using the following chain of programs:
SmeftFR [112] $\to$ FeynArts [137] $\to$ FeynCalc [138]
The new feature of this analysis is the consistent inclusion of $O(1/\Lambda^{4})$ dimension-62 terms. Neglecting such terms e.g. in normalisation factors $Z_{h},Z_{W},\ldots$ between weak eigenbasis and physical fields (see Refs. [118, 112]), leads to error by a factor of ${\cal O}(10)$ in the amplitudes when the dimension-62 terms are relevant. Even more important is the fact that, failure of a consistent truncation in $Z$-normalisation factors here would lead to gauge-fixing parameter dependent results and therefore meaningless $S$-matrix elements. This omission may turn out to be important in setting bounds to Wilson coefficients or predictions for physical observables in SMEFT.
Inspecting the formulae for helicity amplitudes displayed in Appendices A and B one can see that contributions from some dimension-6 operators ($Q_{\varphi\Box},Q_{\varphi D},Q_{\varphi B},Q_{\varphi W},Q_{\varphi WB}$) may grow linearly with $s$. Including dimension-62 terms does not change this conclusion, the new contributions have the form $C^{2}s/(G_{F}\Lambda^{4})$. However, corrections to amplitudes due to some dimension-8 operators ($Q_{\varphi^{4}D^{4}}^{(1)},Q_{\varphi^{4}D^{4}}^{(2)},Q_{\varphi^{4}D^{4}}^{(3)},Q_{B^{2}\varphi^{2}D^{2}}^{(1)},Q_{W^{2}\varphi^{2}D^{2}}^{(1)},Q_{W^{2}\varphi^{2}D^{2}}^{(2)}$) grow faster with energy, like $s^{2}$.
In Figure 5 we illustrate possible enhancement of the $ZZ\to HH$ cross-sections in SMEFT due to effects growing with energy. We present results in terms of the ratio to the SM cross-section, defined as:
$R^{VV}=\frac{\sigma_{SMEFT}(VV\to HH)}{\sigma_{SM}(VV\to HH)}\;,\quad V=(W^{\pm},\,Z)\;.$ (4.7)
We show only the results for the $W^{+}W^{-}\rightarrow HH$ process, as for all operators which contribute to $ZZ$ scattering their $s$-dependence is qualitatively similar to the $W^{+}W^{-}$ case. For all plots and numerical results, we assume one non-vanishing WC at a time and its sign is chosen to maximise the cross-section.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Ratio $R^{WW}$ between angular-integrated cross-sections in the SMEFT and the SM for chosen dimension-6 WCs $C_{\varphi\Box}$, $C_{\varphi}$, $C_{\varphi W}$, including dimension-6 (solid lines) and dimension-62 (dashed lines) terms, and for dimension-8 WCs $C_{\varphi^{8}}$, $C_{\varphi^{6}\Box}$, $C_{\varphi^{4}D^{4}}^{(3)}$, $C_{W^{2}\varphi^{2}D^{2}}^{(1)}$, $C_{W^{2}\varphi^{4}}^{(1)}$. Values of dimension-6 WCs are given on corresponding plots and $\Lambda=1$ TeV is assumed. For dimension-8 WCs, NDA scaling is assumed with numerical values of $g_{*}=\sqrt{2}\frac{\Lambda}{1\text{ TeV}}$ and $\Lambda=4$ TeV.
Plots in Figure 5 are chosen to illustrate several important points. First, one can see that for dimension-6 WCs values close to the upper bounds dictated by fits to experimental data (see Table 3), dimension-62 terms become dominant for $\sqrt{s}\gtrsim 600~\mathrm{GeV}$ (excluding $C_{\varphi}$, for which there are no double insertions at the level of the amplitude, hence the LO and NLO results are identical). Moreover, one can expect significant enhancement of the $VV\rightarrow HH$ cross-section from all the classes of dimension-6 WCs, with specific examples presented in Figure 5: $\varphi^{6}$: $C_{\varphi}$, $\varphi^{4}D^{2}$: $C_{\varphi\Box}$ and $X^{2}\varphi^{2}$: $C_{\varphi W}$. For dimension-8 WCs, the most promising classes are: $\varphi^{8}$: $C_{\varphi^{8}}$, $\varphi^{6}D^{2}$: $C_{\varphi^{6}\Box}$ and $X^{2}\varphi^{4}$: $C_{W^{2}\varphi^{4}}^{(1)}$. The relative magnitude of enhancement from various WCs is primarily dependent on their maximal allowed values given by numerical fits or NDA. The $s$-dependence can, for some operators, further significantly increase the $R^{VV}$ ratio, but one should remember that in the high-$s$ range, the number of produced events is lower and statistical errors of measurements larger, so this tail effect may be more difficult to be observed. However, even if not relevant for the HL-LHC, it might become more of significance for future planned experiments such as FCC-hh.
5 Maximal enhancement of the double Higgs production via VBF in SMEFT
We now turn our attention towards finally answering the question: What is the maximum enhancement in SMEFT of the double Higgs boson production rate through vector boson fusion at the HL-LHC?
The answer requires an analysis of the full process $pp\to hh+\mathrm{jets}$, rather than just the $VV\to hh$ sub-process which we discussed in detail in the previous sections, to analytically understand the most important effects of the higher-dimensional SMEFT operators. To this end, we performed a numerical Monte Carlo analysis of this process for the LHC and its upgrade HL-LHC. The results were obtained using MadGraph 5 - v3_4_1 [139] event generator utilising UFO [140] model files generated with SmeftFR v3.02 [112]. The process of interest is:
p p > h h j j
where p represents the initial protons, h the Higgs bosons, and j the final-state jets. We utilised the VBF-specific cuts on kinematic variables that are based on the ATLAS recommendations (following [130, 93, 101]) on $M_{jj}$ - invariant mass of a pair of jets, $\Delta\eta_{jj}$ - the rapidity separation of jets, $\eta_{j}$ - the rapidity of individual jets, $\Delta p_{Tj}$ - the transverse momentum of jets, and $M_{hh}$ - the invariant mass of the pair of Higgs bosons:
$M_{jj}\geq 700\text{ GeV},\quad\Delta\eta_{jj}\geq 5,\quad|\eta_{j}|\leq 4.5,\quad p_{Tj}\geq 25\text{ GeV},\quad M_{hh}\geq 400\text{ GeV}.$ (5.1)
It should be noted that a more detailed study including Higgs decays to final states, may require more refined and final-state-specific cuts on kinematic variables.
Moreover, we assumed all fermion masses — apart from the top quark — to be equal to zero, and the flavour mixing matrices to be equal to identity. The numerical values of input parameters are in line with the default SmeftFR v3 $G_{F}$” default input scheme and read:
[…]
We run the simulations for the LHC and HL-LHC experiments, in the rest of the paper assuming the energy $\sqrt{s}=13$ TeV and integrated luminosity of $\mathcal{L}=30$ fb-1 for the LHC and $\sqrt{s}=14$ TeV, $\mathcal{L}=3000$ fb-1 for the HL-LHC.
As the maximal values of Wilson coefficients allowed by current experimental data (eventually extrapolated using the NDA assumption) we adopt their numerical values listed in Section 2.2. As some of them depend on the new physics scale $\Lambda$, we employ the EFT validity criterion of Eq. (3.28) to determine its minimal value, limited by the condition that less than $1\%$ of all generated Higgs pairs have invariant mass higher than $\Lambda$ (see Sec. 5.2 for more details).
We start with a study assuming one non-zero WC at a time (with an exception to $C_{\varphi}$ and $C_{\varphi^{8}}$ affecting $\kappa_{\lambda_{3}}$, see discussion in Section 5.1). For each run, we calculate the cross-section and the number-of-event enhancements understood as:
$\Delta\sigma=\frac{\sigma^{SMEFT}}{\sigma^{SM}},\quad\Delta N=N^{SMEFT}-N^{SM}\;.$ (5.3)
Whenever necessary, we further distinguish the contributions from various subclasses of terms, e.g. $\Delta\sigma_{D6},\Delta\sigma_{D6^{2}},\Delta\sigma_{D8}$ denote cross-section enhancement considering, respectively, dimension-6 terms only, dimension-6 and dimension-62 terms or dimension-8 terms only (and similarly for event excesses $\Delta N_{D6},\Delta N_{D6^{2}},\Delta N_{D8}$). We use such a self-explanatory notation in all following Tables and Figures.
We display the results for both positive and negative ends of the allowed WC ranges. Moreover, we present kinematic distributions of Higgs pair invariant mass $M_{hh}$ to illustrate where the BSM effects can be easiest to identify.
5.1 WCs affecting Higgs self-couplings
We start by analysing the enhancement associated with the triple Higgs coupling modifier $\kappa_{\lambda_{3}}$. As discussed in Section 2.2, experimental constraints on $\kappa_{\lambda_{3}}$ can be translated into bounds on two Wilson coefficients that primarily affect this coupling: $C_{\varphi}$ and $C_{\varphi^{8}}$. However, as can be seen from the formulae in Appendices A and B, the dominant contributions from both WCs affect mainly longitudinal scattering amplitudes $\mathcal{M}^{WW(ZZ)}_{00}$ and enter into relevant expression in the same linear combination as into $\kappa_{\lambda_{3}}$ (note also that there are no $C_{\varphi}^{2}$ contributions, distinguishing LO and NLO orders of EFT expansion). Therefore, bounds on di-Higgs production are in this case directly given by current limits on $\kappa_{\lambda_{3}}$ [see Eq. (2.2)], independently of individual values of $C_{\varphi}$ and $C_{\varphi^{8}}$. For illustration, we present them assuming $C_{\varphi}\neq 0$ and $C_{\varphi^{8}}=0$ using limits of Eq.(2.3). The numerical results of the impact of these Wilson coefficients on di-Higgs production are shown in Table 4, while the invariant mass distributions are displayed in Figure 6. As can be seen from Table 4, the enhancement can be substantial, especially in the negative sign case. The distributions in Figure 6 show, consistent with Figure 5, that the most significant excess of events is expected at low energies, with $M_{hh}=\sqrt{s}\lesssim 1000$ GeV.
It should be noted that while in this instance the enhancement can reach factor $\mathcal{O}(20)$, gluon fusion would also cause a similar excess in the dominant production channel, resulting in a much earlier measurement. For this reason, we consider other Wilson coefficients, endemic for VBF double Higgs production, to be more interesting, as discussed in the following sections.
LHC
WC $\Delta\sigma$ $\Delta N$
$\frac{C_{\varphi}}{1\text{ TeV}^{2}}$ $4.7$ 6.0 81
13.8-13.8 22.2 336
HL-LHC
WC $\Delta\sigma$ $\Delta N$
$\frac{C_{\varphi}}{1\text{ TeV}^{2}}$ $4.7$ 5.9 9550
13.8-13.8 21.5 39784
Table 4: Maximal enhancement of the di-Higgs production via VBF for $C_{\varphi}$ modifying $\kappa_{\lambda_{3}}$.
Refer to caption Refer to caption
Figure 6: Higgs pair invariant mass $M_{hh}$ distribution for $C_{\varphi}$ modifying $\kappa_{\lambda_{3}}$.
5.2 WCs affecting Higgs-vector boson interactions
Dimension-6 operators
In the case of dimension-6 Wilson coefficients, we assume numerical values based on the results of global fits cited in Table 3. We use wider, “marginalised” bounds, as we first aim to examine their individual effects, keeping in mind that, in realistic scenarios, they never appear in isolation. We illustrate the magnitude of difference in results when only single insertions of dimension-6 WCs in the amplitudes are allowed (i.e. linear order), with the case in which we allow double insertions (quadratic order). Moreover, as argued in Section 3.4, we do not truncate higher order EFT terms at the level of the cross-section as the consistent and correct choice.
Our results are presented in Table 5 and illustrated in Figure 7 for the $C_{\varphi\Box}$ and $C_{\varphi W}$ WCs, which correspond to the strongest enhancements. These results indicate that the strongest enhancement of di-Higgs VBF production due to dimension-6 Wilson coefficients at the HL-LHC reaches at most 2.7 for $C_{\varphi W}$, and is significantly lower for all remaining WCs. Furthermore, they highlight the importance of NLO (dimension-62) contributions in comparison to LO (dimension-6) ones. For example, in the case of $C_{\varphi\Box}=-2$, the LO contribution is much larger than the NLO one, consistent with the behaviour shown in the plots in Figure 5. This is further illustrated by the Higgs invariant mass histograms in Figure 7. The reason for this difference is the negative interference between the SM, LO, and NLO contributions in the negative-sign case, which can be examined in Eq. (A.2). The histograms in Figure 7 also demonstrate the energy growth of the cross-section, as can be seen in Figure 5. Finally, from these results, one can deduce the EFT validity scale in line with the requirement of Eq. (3.28): $\Lambda\gtrsim 3$ TeV for $C_{\varphi\Box}$ and $\Lambda\gtrsim 4$ TeV for $C_{\varphi W}$.
LHC WC $\frac{C}{1\text{ TeV}^{2}}$ $\Delta\sigma_{D6}$ $\Delta N_{D6}$ $\Delta\sigma_{D6^{2}}$ $\Delta N_{D6^{2}}$ $C_{\varphi\Box}$ 22 1.04 1 1.53 9 2-2 2.41 23 1.41 7 $C_{\varphi D}$ $1.5$ 1.08 1 1.07 1 1.5-1.5 1.00 0 1.02 0 $C_{\varphi W}$ 11 1.74 12 2.49 24 1-1 2.20 19 1.62 10 $C_{\varphi WB}$ 11 1.27 4 1.14 2 1-1 1.23 4 1.41 7 $C_{\varphi B}$ 11 1.39 6 1.69 11 1-1 1.38 6 1.17 3 HL-LHC WC $\frac{C}{1\text{ TeV}^{2}}$ $\Delta\sigma_{D6}$ $\Delta N_{D6}$ $\Delta\sigma_{D6^{2}}$ $\Delta N_{D6^{2}}$ $C_{\varphi\Box}$ 22 1.07 134 1.59 1144 2-2 2.42 2770 1.41 800 $C_{\varphi D}$ $1.5$ 1.07 140 1.07 131 1.5-1.5 1.01 23 1.02 38 $C_{\varphi W}$ 11 1.85 1645 2.72 3340 1-1 2.31 1504 1.67 1303 $C_{\varphi WB}$ 11 1.28 540 1.14 275 1-1 1.25 481 1.44 848 $C_{\varphi B}$ 11 1.40 773 1.72 1396 1-1 1.39 756 1.17 336
Table 5: Maximal enhancement of VBF di-Higgs production process for dimension-6 WCs.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Higgs pair invariant mass $M_{hh}$ distribution for $C_{\varphi\Box}$ and $C_{\varphi W}$.
Dimension-8 operators
In the case of the dimension-8 Wilson coefficients considered in this paper, we assume their numerical values in line with Eq. (2.16), leading to coefficients $C_{\varphi^{4}D^{4}}$, $C_{V^{2}\varphi^{2}D^{2}}$, and $C_{V\varphi^{4}D^{2}}$ scaling with the new physics scale $\Lambda$. To estimate their maximal effect on di-Higgs production, we must first find a minimal scale $\Lambda$ consistent with the validity criterion of Eq. (3.28).
We do that by generating several event samples for $\Lambda=1,2,3,4,5$ TeV, while assigning corresponding numerical values to the Wilson coefficients, assuming their scaling in Eq. (2.16). For each sample, we impose a cut on the Higgs pair invariant mass, assuming $M_{hh}<\sqrt{s_{\text{max}}}=\Lambda$, and define the “minimal” value of $\Lambda$ as the one for which the number of events excluded by this cut is of the order of $\mathcal{O}(1\%)$. We emphasise once more that a more stringent scrutiny of our assumptions would only result in a further reduction of the event excess. This procedure yields maximal absolute values for dimension-8 WCs ranging from $0.13/\mathrm{TeV}^{4}$ to $4/\mathrm{TeV}^{4}$, with EFT validity scale $\Lambda$ generally exceeding 2-4 TeV. Then, at NLO, the most significant enhancements of di-Higgs VBF production at the HL-LHC, compared to Standard Model predictions, are at most a factor of $1.6$ for $C_{W^{2}\varphi^{4}}^{(1)}$ and $1.3$ for $C_{\varphi^{6}\Box}$. Enhancements for other WCs are considerably smaller. These increases are even less pronounced than those from dimension-6 WCs, making them unlikely to produce any observable effects given the expected HL-LHC sensitivity.
5.3 Interference effects of multiple WCs
If one assumes that several WCs do not vanish at the same time, at the NLO level, quadratic contributions from dimension-62 WCs may produce interesting interference effects, rising the possibility of additional di-Higgs production enhancement.
BP1
WC $C_{\varphi}$ $C_{\varphi\Box}$
$\frac{C}{1\text{ TeV}^{2}}$ A 4.7 2
B 4.7 -2
C -13.8 2
D -13.8 -2
BP2
WC $C_{\varphi}$ $C_{\varphi^{8}}$ $C_{\varphi\Box}$ $C_{\varphi^{6}\Box}$
$\frac{C}{1\text{ TeV}^{(2,4)}}$ A 3.8 7.4 2 4
B 3.8 7.4 -2 -4
C -10 -13.6 2 4
D -10 -13.6 -2 -4
BP3
WC $C_{\varphi\Box}$ $C_{\varphi W}$
$\frac{C}{1\text{ TeV}^{(2)}}$ A 2 1
B 2 -1
C -2 1
D -2 -1
BP4
WC $C_{\varphi\Box}$ $C_{\varphi^{6}\Box}$ $C_{\varphi W}$ $C^{(1)}_{W^{2}\varphi^{4}}$
$\frac{C}{1\text{ TeV}^{(2,4)}}$ A 2 4 1 4
B 2 4 -1 -4
C -2 -4 1 4
D -2 -4 -1 -4
Table 6: Four benchmark scenarios used to evaluate possible enhancement of the VBF di-Higgs production process. New physics scale $\Lambda=4$ TeV is assumed. Letters A, B, C and D label sub-scenarios that take into account different signs configuration.
To assess the impact of several selected non-zero Wilson coefficients on VBF double Higgs production, we collect the results for chosen benchmark points defined in Table 6. The relations between WCs in points BP1 and BP2 follow the structure generated in the scalar singlet model discussed in Section 3 (see also e.g. [123]). In such model, the leading contributions arise from the $C_{\varphi}$, $C_{\varphi^{8}}$, $C_{\varphi\Box}$ and $C_{\varphi^{6}\Box}$ WCs. BP1 includes only dimension-6 WCs, whereas BP2 also considers dimension-8 operators [^8] At the amplitude level, dimension-8 WCs add linearly to other terms. However, as we do not truncate higher-order terms in cross sections, also dimension-8 terms may lead to non-trivial interference with other WCs; thus we include them in our benchmark points.. Points BP3 and BP4 include $C_{\varphi\Box}$, $C_{\varphi^{6}\Box}$, $C_{\varphi W}$ and $C_{W^{2}\varphi^{4}}^{(1)}$ WCs (all potentially loop-generated here) and are chosen to illustrate the possible enhancement with $\kappa_{\lambda_{3}}\simeq 1$, since $C_{\varphi}$ and $C_{\varphi^{8}}$ are set to zero.
LHC Scenario $\Delta\sigma_{D6}$ $\Delta N_{D6}$ $\Delta\sigma_{D6^{2}+D8}$ $\Delta N_{D6^{2}+D8}$ BP1 A 3.95 48 6.36 87 B 9.63 140 3.73 44 C 28.17 440 65.64 1046 D 16.97 259 3.98 48 BP2 A 2.90 31 5.17 68 B 7.77 110 5.19 68 C 15.30 231 46.48 736 D 7.63 107 3.16 35 BP3 A 1.96 15 4.73 60 B 2.06 17 2.76 29 C 2.99 32 1.89 14 D 3.77 45 1.64 10 BP4 A 1.96 15 8.00 113 B 2.06 17 5.10 66 C 2.99 32 3.80 45 D 3.77 45 2.90 31 HL-LHC Scenario $\Delta\sigma_{D6}$ $\Delta N_{D6}$ $\Delta\sigma_{D6^{2}+D8}$ $\Delta N_{D6^{2}+D8}$ BP1 A 3.87 5581 6.23 10177 B 9.48 16498 3.69 5230 C 27.69 51934 65.02 124564 D 16.57 30304 3.91 5656 BP2 A 2.87 3634 5.13 8035 B 7.66 12955 5.13 8029 C 15.13 27490 46.08 87724 D 7.53 12700 3.12 4132 BP3 A 2.10 2134 5.21 8197 B 2.21 2356 2.98 3856 C 3.12 4129 1.98 1912 D 3.91 5653 1.66 1282 BP4 A 2.10 2134 8.87 15316 B 2.21 2356 5.62 8983 C 3.12 4129 4.13 6097 D 3.91 5653 3.02 3934
Table 7: Enhancement of VBF di-Higgs production for benchmark scenarios BP1–BP4. $\Delta\sigma_{D6}$ and $\Delta\sigma_{D6^{2}+D8}$ represent cross-section enhancements at linear and full quadratic order (i.e., including in the amplitude dimension-6 terms only or also dimension-62 and dimension-8 contributions), respectively. $\Delta N_{D6}$ and $\Delta N_{D6^{2}+D8}$ indicate the respective event yield enhancements.
The results of Monte Carlo simulations are summarised in Table 7 and in the histograms of Figure 8. One can draw the following conclusions:
  • BP1 - The enhancement is primarily driven by modifications of $\kappa_{\lambda_{3}}$, resulting in scenario C in maximal increase of $\mathcal{O}(30)$ at linear order and up to $\mathcal{O}(65)$ at quadratic order, for both the LHC and HL-LHC. With luck, such a significant effect could be observed even before the HL-LHC era. Quadratic effects are large and have to be taken into account for viable cross-section estimates, especially that at the NLO order they can lead to both significant enhancement or suppression of the final result, depending on relative WC signs.
  • BP2 - The conclusions mirror those for BP1, with the main difference arising from the inclusion of $C_{\varphi^{8}}$ in the $\kappa_{\lambda_{3}}$ modification, along with additional contributions from $C_{\varphi^{6}\Box}$.
  • BP3 - In the absence of contributions to $\kappa_{\lambda_{3}}$, the derivative WCs $C_{\varphi\Box}$ and $C_{\varphi W}$ can only produce smaller enhancement, reaching $\mathcal{O}(5)$ in Scenario A. Again, quadratic effects are large and can lead to both constructive or destructive interference between WCs.
  • BP4 - The conclusions are consistent with those for BP3, with the enhancement of the cross-section reaching $\mathcal{O}(9)$ after including also WCs of dimension-8 operators $C_{\varphi^{6}\Box}$ and $C_{W^{2}\varphi^{4}}^{(1)}$.
These results highlight the importance of including quadratic contributions in the analysis, both when calculating total cross sections and in finding the shape of invariant mass distribution (as illustrated in Figure 8). Although operators related to $\kappa_{\lambda_{3}}$ dominate the enhancement, interference between $C_{\varphi}$, $C_{\varphi^{8}}$ and other WCs can lead to large effects, potentially comparable to those in the gluon fusion channel. The important role is played here by the relative signs of various WCs. Even without tree-level triple Higgs boson coupling modification, i.e. when $\kappa_{\lambda_{3}}=1$, enhancement from derivative WCs can reach up to $\mathcal{O}(10)$. This conclusion should be interpreted with caution, as we do not include the full set of dimension-8 operators; the result may therefore depend on the choice of operator basis.
We should also comment that our analysis is preliminary and thus does not specify more refined experimental cuts. In case of momentum-dependent Higgs-vector boson couplings, which are particularly interesting as they are weakly or not at all probed by the ggF channel, even a simple cut on invariant mass of the Higgs pair may further improve a ratio of signal from momentum dependent operators from the SM background. In Table 8 we show the results obtained assuming $M_{hh}\geq 1$ TeV. As one can see, comparing to numbers in Tables 5 and 7, such a cut can increase the signal to SM prediction ratio by additional factors of up to almost 5, while still leaving the total number of events large enough to be observable and statistically significant.
HL-LHC
WC $\frac{C}{1\text{ TeV}^{2,4}}$ $\Delta\sigma_{D6}$ $\Delta N_{D6}$ $\Delta\sigma_{D6^{2}+D8}$ $\Delta N_{D6^{2}+D8}$
$C_{\varphi\Box}$ 22 2.62 332 4.96 810
$C_{\varphi W}$ 11 6.35 1095 11.12 2070
$C_{\varphi^{6}\Box}$ 4-4 - - 1.43 88
$C_{W^{2}\varphi^{4}}^{(1)}$ 4-4 - - 3.80 573
BP3 scenario A 7.92 1415 23.09 4517
BP4 scenario A 7.92 1415 41.86 8357
Table 8: Maximal enhancement of VBF di-Higgs production for dimension-6 Wilson coefficients (WCs) obtained after applying $M_{hh}\geq 1$ TeV cut.
The presented results should be put into a context with expected experimental accuracy. Following the recent ATLAS report [141] on the projected HL-LHC sensitivity to double Higgs boson production, we quote it for the VBF-induced double Higgs production as
$\mu_{\text{VBF}-hh}\equiv\Delta\sigma\gtrsim 10\;.$ (5.4)
This value is significantly above the enhancements predicted in the previous sections, where only a single Wilson coefficient was considered at a time, leading to at most $\Delta\sigma\lesssim 3$ and limited prospects for observing VBF di-Higgs boson production at the HL-LHC within the regime of EFT validity.
However, as demonstrated in this section, in more elaborated multiple WC scenarios, or using cuts on invariant mass of the Higgs pair, the enhancement can become large enough to make the VBF channel detectable in future, in parallel to the ggF channel. Measuring (or putting upper limits) on both of them is valuable, as the sets of SMEFT WCs which they probe overlap only partially. The results shown highlight also the importance of considering the quadratic terms in the EFT expansion and interference-driven effects between multiple Wilson coefficients. Such effects can produce sizeable cross-section enhancement in the VBF channel, even in cases where the ggF channel remains unchanged.
Figure 8: Higgs pair invariant mass $M_{hh}$ distribution for chosen BPs from Table 6 for HL-LHC.
6 Conclusions
Our study explores the influence of higher-dimensional operators of the Standard Model Effective Field Theory on the double Higgs boson production process via Vector Boson Fusion. Although the VBF channel is subdominant compared to gluon fusion, it is characterised by a unique experimental signature and provides sensitivity to a different spectrum of BSM interactions.
We incorporated all bosonic CP-conserving dimension-6 operators and a chosen representative subset of dimension-8 operators, arranged in Table 2. These operators may significantly impact the tree-level Higgs bosons self-interactions, and its interactions with $W$ and $Z$ gauge bosons. [^9] In SMEFT there are also $VV^{\prime}h$ and $VV^{\prime}hh$ vertices with $V,V^{\prime}={\gamma,Z}$ that are driven by $C_{\varphi W},C_{\varphi B},C_{\varphi WB}$ WCs. These are, of course, included in our numerical results but not shown analytically in the Appendices. For the dimension-8 operators, we believe that our selection is adequate for capturing the overarching conclusions regarding the maximum possible enhancement of the VBF double Higgs production process within SMEFT. Our choice encompasses all main classes of dimension-8 operators corresponding to possible configurations of energy growth in the amplitudes. The complete set of such operators is very large and would obscure, rather than clarify, the conclusions and discussion presented here. Moreover, since the specific set of operators generated depends on the underlying UV model, a broader selection would not necessarily be more valid or useful. Our aim in this work is to illustrate representative patterns of discussed behavior, which can then be straightforwardly extrapolated to more realistic scenarios.
We determined realistic numerical values for the Wilson coefficients associated with the chosen SMEFT operators in Section 2.2. This involved either leveraging existing bounds (as detailed in Table 3), that are derived from numerical fits to experimental data, or extrapolating such bounds to currently unconstrained WCs by applying Naive Dimension Analysis assumptions. To further illustrate these concepts, we employed examples of singlet and triplet scalar extensions of the SM in Section 3. This enabled us to demonstrate when and how such Wilson coefficient values could be realised after the decoupling of complete UV-models. We utilised a natural matching technique, deployed directly in the SMEFT mass basis, and leveraged the Goldstone Boson Equivalence Theorem. This allowed us to directly map UV-model parameters to the relevant SMEFT Wilson coefficients, as evident from the high energy analytical forms of specific amplitudes (provided in the Appendices A and B). Concurrently, we also examined the convergence criteria for the Effective Field Theory expansion using these same SM scalar extensions.
Subsequently, in Sections 4 and 5, we conducted an in-depth analysis of the $VV\to hh$ and the full $pp\to hhjj$ cross-sections, respectively. Our analysis included the full suite of next-to-leading order bosonic contributions (in the EFT expansion), specifically dimension-6, dimension-62, and dimension-8 terms, which contribute to the tree-level amplitudes. Through this analysis, we pinpointed which WCs exert the most significant impact and how these contributions vary with the invariant mass of the Higgs boson pair in the final state. The key findings and implications of our work, arranged in Tables 4-8 and Figures 6-8 are summarised as follows:
  1. 1.
    Non-derivative operators ($C_{\varphi},\,C_{\varphi^{8}}$): these WCs contribute to the trilinear Higgs coupling and are by far the most significant in potential enhancement of the VBF di-Higgs cross-section, by a factor of at most $\mathcal{O}(20)$ (see Table 4). Nevertheless, we also note that this kind of modification would also drastically change the ggF channel, which could result in its much earlier detection.
  2. 2.
    Derivative operators (e.g. $C_{\varphi\Box},\,C_{\varphi W}$): the VBF di-Higgs production cross-section exhibits only modest enhancements, typically at most a factor of $\mathcal{O}(2-3)$ for dimension-6 WCs, as presented in Table 5. Even smaller improvement can be seen in dimension-8 WCs, where the increase is no more than of order $\mathcal{O}(1.5)$. Consequently, these effects are difficult to detect at the LHC and continue to be marginal with the High-Luminosity LHC (HL-LHC) upgrade. Nevertheless, we noticed that the inclusion of dimension-62 terms in the amplitude, may alter the predicted enhancement by $\approx 50\%$ baring cancellations with the full set of dimension-8 operators.
  3. 3.
    In scenarios where multiple WCs contribute constructively, particularly those involving derivative operators (e.g., $C_{\varphi\Box}$, $C_{\varphi W}$), the cross-section can be enhanced by up to an order of magnitude. Such enhancements are most pronounced in kinematic regions with high invariant mass of the Higgs pair ($M_{hh}\gtrsim$ 1 TeV), where the energy-dependent terms dominate – see Tables 6 and 8. This underscores the significance of incorporating contributions from dimension-62 and dimension-8 operators, as their interference can substantially alter predictions.
In summary, while the VBF di-Higgs channel alone may not provide definitive evidence of BSM physics at the HL-LHC, it serves as an important complementary probe, especially for derivative interactions and more realistic, multi-WC scenarios. Our results highlight the importance of including higher-order EFT contributions, and emphasise the need for a careful assessment of EFT validity, motivating further experimental and theoretical efforts to explore this process in greater detail.
Acknowledgements
The authors acknowledge support from the COMETA COST Action CA22130. The work of JR was supported by the Polish National Science Center under the research grant DEC2019/35/B/ST2/02008. MR acknowledges support from the University of Padua under the 2023 STARS Grants@Unipd programme (Acronym and title of the project: HiggsPairs – Precise Theoretical Predictions for Higgs pair production at the LHC). We would like to thank Stefan Pokorski for numerous fruitful discussions.
Appendices: Dominant contributions to helicity amplitudes
In the following two sections, we present the helicity amplitudes for the $WW\rightarrow HH$ and $ZZ\rightarrow HH$ processes, including SMEFT contributions from the dimension-6 and dimension-8 operators collected in Table 2. For each amplitude, we display only the leading order terms in the high energy limit ($M_{V}^{2}/s\rightarrow 0$), including interference terms between different Wilson coefficients. We start with the $WW\rightarrow HH$ process in section A and present the Standard Model contribution in Eq. (A.1), the dimension-6 and dimension-62 contributions in Eq. (A.2), and the dimension-8 contributions in Eqs. (A.3) and (A.4). Similarly, in section B, for the $ZZ\rightarrow HH$ process, we present the SM contribution in Eq. (B.1), the dimension-6 and dimension-62 contributions in Eq. (B.2), and the dimension-8 contributions in Eqs. (B.3) and (B.4).
Appendix A $W^{+}W^{-}\rightarrow HH$ helicity amplitude
A.1 SM
$\mathcal{M}^{WW,\,SM}_{00}$ $=\sqrt{2}G_{F}M_{H}^{2}\mathopen{}\mathclose{{\left[1+2\mathopen{}\mathclose{{\left(1-\frac{4}{\sin^{2}\theta}}}\right)\frac{M_{W}^{2}}{M_{H}^{2}}}}\right]\;,$ (A.1)
$\mathcal{M}^{WW,\,SM}_{\pm\pm}$ $=\mathcal{O}\mathopen{}\mathclose{{\left(\frac{1}{s}}}\right)\;,$
$\mathcal{M}^{WW,\,SM}_{\pm\mp}$ $=2\sqrt{2}G_{F}M_{W}^{2}\;,$
$\mathcal{M}^{WW,\,SM}_{0\pm}$ $=\mathcal{M}^{WW,\,SM}_{\pm 0}=\mathcal{O}\mathopen{}\mathclose{{\left(\frac{1}{\sqrt{s}}}}\right)\,.$
A.2 Dimension-6 SMEFT
$\mathcal{M}^{WW,\,D6}_{00}=$ $-3\sqrt{2}C_{\varphi}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)-20C_{\varphi W}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)+\frac{1}{2}(4C_{\varphi\Box}-C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\hskip 85.35826pt\textbf{dim-6}$
$+\frac{1}{2\sqrt{2}}(4C_{\varphi\Box}-C_{\varphi D})^{2}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)-3C_{\varphi}(4C_{\varphi\Box}-C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)^{2}\;,\hskip 36.98866pt\textbf{(dim-6)}^{2}$
$\mathcal{M}^{WW,\,D6}_{\pm\pm}=$ $-(4C_{\varphi\Box}-C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)+2C_{\varphi W}\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)$
$+\sqrt{2}C_{\varphi W}(4C_{\varphi\Box}-C_{\varphi D}+6C_{\varphi W})\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)-6C_{\varphi}C_{\varphi W}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)^{2}\;,$ (A.2)
$\mathcal{M}^{WW,\,D6}_{\pm\mp}=$ $+(4C_{\varphi\Box}-C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)$
$+2\sqrt{2}C_{\varphi W}^{2}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$
$\mathcal{M}^{WW,\,D6}_{0\pm}=$ $-4\sqrt{2}\cot\theta C_{\varphi W}\mathopen{}\mathclose{{\left(\frac{\sqrt{s}M_{W}}{\Lambda^{2}}}}\right)$
$-2C_{\varphi W}\mathopen{}\mathclose{{\left(4C_{\varphi W}+4C_{\varphi\Box}-C_{\varphi D}}}\right)\cot\theta\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{\sqrt{s}M_{W}}{\Lambda^{2}}}}\right)\;.$
A.3 Dimension-8 SMEFT
For clarity, we divide the expressions for the dimension-8 terms into 2 classes.
A.3.1 $\varphi^{8}$, $\varphi^{6}D^{2}$, $\varphi^{4}D^{4}$ operator classes
$\mathcal{M}^{WW,\,D8}_{00}=$ $-6C_{\varphi^{8}}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}^{2}\Lambda^{4}}}}\right)+\frac{\sqrt{2}}{4}\mathopen{}\mathclose{{\left(8C_{\varphi^{6}\Box}-C_{\varphi^{6}D^{2}}}}\right)\frac{s}{\Lambda^{2}}\frac{1}{G_{F}\Lambda^{2}}$
$-\frac{1}{8}\,\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(C_{\varphi^{4}D^{4}}^{(1)}+C_{\varphi^{4}D^{4}}^{(2)}+4\,C_{\varphi^{4}D^{4}}^{(3)}}}\right)+\mathopen{}\mathclose{{\left(C_{\varphi^{4}D^{4}}^{(1)}+C_{\varphi^{4}D^{4}}^{(2)}}}\right)\cos^{2}\theta}}\right]\,\,\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)^{2}\;,$
$\mathcal{M}^{WW,\,D8}_{\pm\pm}=$ $-\frac{1}{\sqrt{2}}\,\mathopen{}\mathclose{{\left(8\,C_{\varphi^{6}\Box}-C_{\varphi^{6}D^{2}}}}\right)\,\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)\,\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)$
$+\frac{1}{4}\mathopen{}\mathclose{{\left[4\,C_{\varphi^{4}D^{4}}^{(3)}+\mathopen{}\mathclose{{\left(C_{\varphi^{4}D^{4}}^{(1)}+C_{\varphi^{4}D^{4}}^{(2)}}}\right)\,\sin^{2}\theta}}\right]\,\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)\,\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$ (A.3)
$\mathcal{M}^{WW,\,D8}_{\pm\mp}=$ $+\frac{\sqrt{2}}{4}\mathopen{}\mathclose{{\left(8C_{\varphi^{6}\Box}-C_{\varphi^{6}D^{2}}}}\right)\frac{M_{Z}^{2}}{\Lambda^{2}}\frac{1}{G_{F}\Lambda^{2}}+\frac{1}{4}\mathopen{}\mathclose{{\left(C_{\varphi^{4}D^{4}}^{(1)}+C_{\varphi^{4}D^{4}}^{(2)}}}\right)\sin^{2}\theta\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$
$\mathcal{M}^{WW,\,D8}_{0\pm}=$ $+\frac{1}{8\sqrt{2}}\mathopen{}\mathclose{{\left(C_{\varphi^{4}D^{4}}^{(1)}+C_{\varphi^{4}D^{4}}^{(2)}}}\right)\sin{2\theta}\mathopen{}\mathclose{{\left(\frac{s^{3/2}M_{W}}{\Lambda^{4}}}}\right)\;.$
A.3.2 $X^{2}\varphi^{4}$, $X^{2}\varphi^{2}D^{2}$, $X\varphi^{4}D^{2}$ operator classes
$\mathcal{M}^{WW,\,D8}_{00}=$ $+\mathopen{}\mathclose{{\left[\frac{1}{2}C_{W^{2}\varphi^{2}D^{2}}^{(1)}(\sin^{2}\theta+1)+2C_{W^{2}\varphi^{2}D^{2}}^{(2)}}}\right]\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)$
$-14\sqrt{2}C_{W^{2}\varphi^{4}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)+\frac{5i}{2^{5/4}}C_{W\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$
$\mathcal{M}^{WW,\,D8}_{\pm\pm}=$ $-\mathopen{}\mathclose{{\left(\frac{1}{4}C_{W^{2}\varphi^{2}D^{2}}^{(1)}+C_{W^{2}\varphi^{2}D^{2}}^{(2)}}}\right)\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)$
$+3\sqrt{2}C_{W^{2}\varphi^{4}}^{(1)}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)-\frac{3i}{2^{5/4}}C_{W\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$ (A.4)
$\mathcal{M}^{WW,\,D8}_{\pm\mp}=$ $-\frac{1}{8}\sin^{2}\theta C_{W^{2}\varphi^{2}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)+2^{3/4}C_{W\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)\;,$
$\mathcal{M}^{WW,\,D8}_{0\pm}=$ $+2^{1/4}iC_{W\varphi^{4}D^{2}}^{(1)}\cot\theta\mathopen{}\mathclose{{\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{M_{W}\sqrt{s}}{\Lambda^{2}}}}\right)$
$-4C_{W^{2}\varphi^{4}}^{(1)}\cot{\theta}\mathopen{}\mathclose{{\left(\frac{s^{1/2}M_{W}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)-\frac{1}{8\sqrt{2}}C_{W^{2}\varphi^{2}D^{2}}^{(1)}\sin{2\theta}\mathopen{}\mathclose{{\left(\frac{s^{3/2}M_{W}}{\Lambda^{4}}}}\right)\;.$
Appendix B $ZZ\rightarrow HH$ helicity amplitude
B.1 SM
$\mathcal{M}^{ZZ,\,SM}_{00}$ $=\sqrt{2}G_{F}M_{H}^{2}\mathopen{}\mathclose{{\left[1+2\mathopen{}\mathclose{{\left(1-\frac{4}{\sin^{2}\theta}}}\right)\frac{M_{Z}^{2}}{M_{H}^{2}}}}\right]\;,$ (B.1)
$\mathcal{M}^{ZZ,\,SM}_{\pm\pm}$ $=\mathcal{O}\mathopen{}\mathclose{{\left(\frac{1}{s}}}\right)\;,$
$\mathcal{M}^{ZZ,\,SM}_{\pm\mp}$ $=2\sqrt{2}G_{F}M_{Z}^{2}\;,$
$\mathcal{M}^{ZZ,\,SM}_{0\pm}$ $=\mathcal{M}^{ZZ,\,SM}_{\pm 0}=\mathcal{O}\mathopen{}\mathclose{{\left(\frac{1}{\sqrt{s}}}}\right)\;.$
B.2 Dimension-6 SMEFT
$\mathcal{M}^{ZZ,\,D6}_{00}=$ $-3\sqrt{2}C_{\varphi}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)+(2C_{\varphi\Box}+C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\hskip 190.63338pt\textbf{dim-6}$
$-20C_{\varphi W}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)+20C_{\varphi B}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}-M_{Z}^{2}}{\Lambda^{2}}}}\right)-20C_{\varphi WB}\mathopen{}\mathclose{{\left(\frac{M_{W}\sqrt{M_{Z}^{2}-M_{W}^{2}}}{\Lambda^{2}}}}\right)$
$+\sqrt{2}(4C_{\varphi\Box}^{2}-\frac{3}{4}C_{\varphi D}^{2})\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)-\frac{3}{2}C_{\varphi}(8C_{\varphi\Box}-C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)^{2}\;,\hskip 48.36958pt\textbf{(dim-6)}^{2}$
$\mathcal{M}^{ZZ,\,D6}_{\pm\pm}=$ $-2(2C_{\varphi\Box}+C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)$
$+2C_{\varphi W}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)+2C_{\varphi B}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)+2C_{\varphi WB}\frac{M_{W}}{M_{Z}}\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)$
$+6\sqrt{2}C_{\varphi W}^{2}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)$
$+6\sqrt{2}C_{\varphi B}^{2}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)+2\sqrt{2}C_{\varphi WB}^{2}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)$
$-6C_{\varphi}\mathopen{}\mathclose{{\left[C_{\varphi W}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)+C_{\varphi B}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)+C_{\varphi WB}\mathopen{}\mathclose{{\left(\frac{M_{W}}{M_{Z}}}}\right)\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}}\right]\mathopen{}\mathclose{{\left(\frac{1}{\Lambda^{2}G_{F}}}}\right)^{2}$
$+\Biggl[-\frac{1}{\sqrt{2}}C_{\varphi W}C_{\varphi D}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)+4\sqrt{2}C_{\varphi W}C_{\varphi\Box}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)$
$-\sqrt{2}C_{\varphi B}C_{\varphi D}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{2M_{Z}^{2}}}}\right)+4\sqrt{2}C_{\varphi B}C_{\varphi\Box}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)$ (B.2)
$-\frac{3\sqrt{2}}{4}C_{\varphi WB}C_{\varphi D}\frac{M_{W}}{\sqrt{M_{Z}^{2}-M_{W}^{2}}}\mathopen{}\mathclose{{\left(1-\frac{2M_{W}^{2}}{3M_{Z}^{2}}}}\right)+4\sqrt{2}C_{\varphi WB}C_{\varphi\Box}\frac{M_{W}}{M_{Z}}\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}$
$+5\sqrt{2}C_{\varphi B}C_{\varphi WB}\mathopen{}\mathclose{{\left(\frac{M_{W}}{M_{Z}}}}\right)\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}+7\sqrt{2}C_{\varphi B}C_{\varphi WB}\mathopen{}\mathclose{{\left(\frac{M_{W}}{M_{Z}}}}\right)\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}\Biggr]\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$
$\mathcal{M}^{ZZ,\,D6}_{\pm\mp}=$ $+(4C_{\varphi\Box}+C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)$
$+\Biggl[4\sqrt{2}C_{\varphi W}^{2}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)+4\sqrt{2}C_{\varphi B}^{2}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)+\sqrt{2}C_{\varphi WB}^{2}$
$+2\sqrt{2}C_{\varphi WB}(C_{\varphi W}+C_{\varphi B)}\mathopen{}\mathclose{{\left(\frac{M_{W}}{M_{Z}}}}\right)\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}\Biggr]\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$
$\mathcal{M}^{ZZ,\,D6}_{0\pm}=$ $-4\sqrt{2}\cot\theta C_{\varphi W}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{\sqrt{s}M_{Z}}{\Lambda^{2}}}}\right)$
$-4\sqrt{2}\cot\theta C_{\varphi B}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{\sqrt{s}M_{Z}}{\Lambda^{2}}}}\right)-4\sqrt{2}\cot\theta C_{\varphi WB}\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}\mathopen{}\mathclose{{\left(\frac{\sqrt{s}M_{W}}{\Lambda^{2}}}}\right)$
$+\Biggl[-16\cot\theta C_{\varphi W}^{2}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)$
$-16\cot\theta C_{\varphi B}^{2}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{\sqrt{s}M_{Z}}{\Lambda^{2}}}}\right)-8\cot\theta C_{\varphi WB}^{2}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)$
$-2\cot\theta C_{\varphi W}(4C_{\varphi\Box}+C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)-2\cot\theta C_{\varphi B}\mathopen{}\mathclose{{\left(4C_{\varphi\Box}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)+C_{\varphi D}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)}}\right)$
$-8\cot\theta C_{\varphi WB}C_{\varphi\Box}\mathopen{}\mathclose{{\left(\frac{M_{W}}{M_{Z}}}}\right)\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}-\cot\theta C_{\varphi WB}C_{\varphi D}\mathopen{}\mathclose{{\left(\frac{M_{W}}{M_{Z}}}}\right)\mathopen{}\mathclose{{\left(1-2\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)^{-1/2}$
$-4C_{\varphi WB}(3C_{\varphi W}+C_{\varphi B)}\mathopen{}\mathclose{{\left(\frac{M_{W}}{M_{Z}}}}\right)\sqrt{1-\frac{M_{W}^{2}}{M_{Z}^{2}}}\Biggr]\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{\sqrt{s}M_{Z}}{\Lambda^{2}}}}\right)\;.$
B.3 Dimension-8 SMEFT
B.3.1 $\varphi^{8}$, $\varphi^{6}D^{2}$, $\varphi^{4}D^{4}$ operator classes
$\mathcal{M}^{ZZ,\,D8}_{00}=$ $-6C_{\varphi^{8}}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}^{2}\Lambda^{4}}}}\right)+\sqrt{2}\mathopen{}\mathclose{{\left(2C_{\varphi^{6}\Box}+C_{\varphi^{6}D^{2}}}}\right)\frac{s}{\Lambda^{2}}\frac{1}{G_{F}\Lambda^{2}}$
$-\frac{1}{2}\,\mathopen{}\mathclose{{\left(C_{\varphi^{4}D^{4}}^{(1)}-\frac{1}{2}C_{\varphi^{4}D^{4}}^{(2)}\sin^{2}\theta+\,C_{\varphi^{4}D^{4}}^{(3)}}}\right)\,\,\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)^{2}\;,$
$\mathcal{M}^{ZZ,\,D8}_{\pm\pm}=$ $-2\sqrt{2}\mathopen{}\mathclose{{\left(2C_{\varphi^{6}\Box}+C_{\varphi^{6}D^{2}}}}\right)\frac{M_{Z}^{2}}{\Lambda^{2}}\frac{1}{G_{F}\Lambda^{2}}$
$+\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(C_{\varphi^{4}D^{4}}^{(1)}-\frac{1}{2}C_{\varphi^{4}D^{4}}^{(2)}+C_{\varphi^{4}D^{4}}^{(3)}}}\right)-\frac{1}{2}C_{\varphi^{4}D^{4}}^{(2)}\cos^{2}\theta}}\right]\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$ (B.3)
$\mathcal{M}^{ZZ,\,D8}_{\pm\mp}=$ $+\sqrt{2}\mathopen{}\mathclose{{\left(2C_{\varphi^{6}\Box}+\frac{3}{4}C_{\varphi^{6}D^{2}}}}\right)\frac{M_{Z}^{2}}{\Lambda^{2}}\frac{1}{G_{F}\Lambda^{2}}+\frac{1}{2}C_{\varphi^{4}D^{4}}^{(2)}\sin^{2}\theta\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$
$\mathcal{M}^{ZZ,\,D8}_{0\pm}=$ $+\frac{1}{4\sqrt{2}}C_{\varphi^{4}D^{4}}^{(2)}\sin{2\theta}\mathopen{}\mathclose{{\left(\frac{s^{3/2}M_{Z}}{\Lambda^{4}}}}\right)\;.$
B.3.2 $X^{2}\varphi^{4}$, $X^{2}\varphi^{2}D^{2}$, $X\varphi^{4}D^{2}$ operator classes
$\mathcal{M}^{ZZ,\,D8}_{00}=$ $+14\sqrt{2}C_{B^{2}\varphi^{4}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}-M_{Z}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)+14\sqrt{2}C_{W^{2}\varphi^{4}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)$
$+\mathopen{}\mathclose{{\left[\frac{1}{2}C_{B^{2}\varphi^{2}D^{2}}^{(1)}(\sin^{2}\theta+1)+2C_{B^{2}\varphi^{2}D^{2}}^{(2)}}}\right]\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)$
$+\mathopen{}\mathclose{{\left[\frac{1}{2}C_{W^{2}\varphi^{2}D^{2}}^{(1)}(\cos^{2}\theta+1)+2C_{W^{2}\varphi^{2}D^{2}}^{(2)}}}\right]\frac{M_{W}^{2}}{M_{Z}^{2}}\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)$
$-\frac{5i}{2^{5/4}}C_{B\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{\sqrt{M_{Z}^{2}-M_{W}^{2}}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)+\frac{5i}{2^{5/4}}C_{W\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$
$\mathcal{M}^{ZZ,\,D8}_{\pm\pm}=$ $+3\sqrt{2}C_{B^{2}\varphi^{4}}^{(1)}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)+3\sqrt{2}C_{W^{2}\varphi^{4}}^{(1)}\frac{M_{W}^{2}}{M_{Z}^{2}}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)$
$-\mathopen{}\mathclose{{\left[\frac{1}{4}C_{B^{2}\varphi^{2}D^{2}}^{(1)}+C_{B^{2}\varphi^{2}D^{2}}^{(2)}}}\right]\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)-\mathopen{}\mathclose{{\left[\frac{1}{4}C_{W^{2}\varphi^{2}D^{2}}^{(1)}+C_{W^{2}\varphi^{2}D^{2}}^{(2)}}}\right]\frac{M_{W}^{2}}{M_{Z}^{2}}\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)$
$-\frac{3i}{2^{5/4}}C_{B\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{\sqrt{M_{Z}^{2}-M_{W}^{2}}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)-\frac{3i}{2^{5/4}}C_{W\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$ (B.4)
$\mathcal{M}^{ZZ,\,D8}_{\pm\mp}=$ $-\frac{1}{8}C_{B^{2}\varphi^{2}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\sin^{2}\theta\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)-\frac{1}{8}C_{W^{2}\varphi^{2}D^{2}}^{(1)}\frac{M_{W}^{2}}{M_{Z}^{2}}\sin^{2}\theta\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)$
$+2^{3/4}iC_{B\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{\sqrt{M_{Z}^{2}-M_{W}^{2}}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)+2^{3/4}iC_{W\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)\;,$
$\mathcal{M}^{ZZ,\,D8}_{0\pm}=$ $-4C_{B^{2}\varphi^{4}}^{(1)}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\cot{\theta}\mathopen{}\mathclose{{\left(\frac{s^{1/2}M_{Z}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)-4C_{W^{2}\varphi^{4}}^{(1)}\frac{M_{W}^{2}}{M_{Z}^{2}}\cot{\theta}\mathopen{}\mathclose{{\left(\frac{s^{1/2}M_{Z}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)$
$-\frac{1}{8\sqrt{2}}C_{B^{2}\varphi^{2}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\sin{2\theta}\mathopen{}\mathclose{{\left(\frac{s^{3/2}M_{Z}}{\Lambda^{4}}}}\right)-\frac{1}{8\sqrt{2}}C_{W^{2}\varphi^{2}D^{2}}^{(1)}\frac{M_{W}^{2}}{M_{Z}^{2}}\sin{2\theta}\mathopen{}\mathclose{{\left(\frac{s^{3/2}M_{Z}}{\Lambda^{4}}}}\right)$
$+2^{1/4}iC_{B\varphi^{4}D^{2}}^{(1)}\cot\theta\mathopen{}\mathclose{{\left(\frac{\sqrt{M_{Z}^{2}-M_{W}^{2}}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{M_{Z}\sqrt{s}}{\Lambda^{2}}}}\right)+2^{1/4}iC_{W\varphi^{4}D^{2}}^{(1)}\cot\theta\mathopen{}\mathclose{{\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{M_{Z}\sqrt{s}}{\Lambda^{2}}}}\right)\;.$